### v1

# Early Directional Convergence in Deep Homogeneous Neural Networks for Small Initializations

Abstract
This paper studies the gradient flow dynamics that arise when training deep homogeneous neural networks, starting with small initializations. The present work considers neural networks that are assumed to have locally Lipschitz gradients and an order of homogeneity strictly greater than two. This paper demonstrates that for sufficiently small initializations, during the early stages of training, the weights of the neural network remain small in norm and approximately converge in direction along the Karush-Kuhn-Tucker (KKT) points of the neural correlation function introduced in [1]. Additionally, for square loss and under a separability assumption on the weights of neural networks, a similar directional convergence of gradient flow dynamics is shown near certain saddle points of the loss function.
1 Introduction
Neural networks have achieved remarkable success across various tasks, yet the precise mechanism driving this success remains theoretically elusive. The training of neural networks involve optimizing a non-convex loss function, where the training algorithm typically is a first-order methods such as gradient descent or its variants. A particularly puzzling aspect is how these training algorithms succeed in finding a solution with good generalization capabilities despite the non-convexity of the loss landscape. In addition to the choice of the training algorithm, the choice of initialization in these algorithms play a crucial role in determining the neural network performance. Indeed, recent works have made increasingly clear the benefit of small initializations, revealing that neural networks trained using (stochastic) gradient descent with small initializations exhibit feature learning [2] and also generalize better for various tasks [3, 4, 5]; see Section 2 for more details into the impact of initialization scale. However, for small initializations, the training dynamics of neural networks is extremely non-linear and not well understood so far. Our focus in this paper is on understanding the effect of small initialization on the training dynamics of neural networks. In pursuit of a deeper understanding of the training mechanism for small initializations, researchers have uncovered the phenomenon of directional convergence in the neural network weights during early phases of training [6, 7]. In [6], authors study the gradient flow dynamics of training two-layer Rectified Linear Unit (ReLU) neural networks. They demonstrate that in the early stages of training, the weights of two-layer ReLU neural networks converge in direction while their norms remain small. This phenomenon is referred to as early directional convergence. In a recent work [1], early directional convergence was further established for two-homogeneous [^1] A neural network $\mathcal{H}$ is defined to be $L$-(positively) homogeneous, if its output $\mathcal{H}(\mathbf{x};\mathbf{w})$ satisfies $\mathcal{H}(\mathbf{x};c\mathbf{w})=c^{L}\mathcal{H}(\mathbf{x};\mathbf{w}),\forall c\geq 0,$ where $\mathbf{x}$ is the input, $\mathbf{w}$ is a vector containing all the weights, and $L$ is referred to as the order of homogeneity. neural networks, illustrating the importance of homogeneity for such types of phenomenon (as opposed to the specific choice of activation function or architecture). Despite focusing solely on the early phases of training, these works demonstrate that interesting behaviors begin to emerge very early in the training dynamics. Moreover, these insights in the initial stages of training have been useful towards a comprehensive understanding of the entire training dynamics in some special cases [8, 9, 10]. So far, all the works on the early stages of training dynamics of neural networks for small initialization have only considered two-layer neural networks (or more broadly two-homogeneous). In contrast, our understanding for deep neural networks is much more limited. In this work, we take a first step towards this goal by focusing on deep homogeneous neural networks (with locally Lipschitz gradients) and establish early directional convergence for them. To our knowledge, there has been no prior work that rigorously demonstrates early directional converge for deeper neural networks. To describe the phenomenon of early directional convergence, [1] introduced the notion of Neural Correlation Function (NCF). For a given neural network and a vector, the NCF quantifies the correlation between the output of neural network and the vector; we provide a more precise definition in Section 4. They show that in the early stages of training, the weights of the neural network remain small and converge in direction to a KKT point of the constrained NCF, an optimization problem that maximizes the NCF on unit sphere. Furthermore, they also establish similar directional convergence near certain saddle points as well. We emphasize that their results are specific to two-homogeneous neural networks. The present work studies the phenomenon of early directional convergence for deep homogeneous neural networks. We particularly consider $L$-homogeneous neural networks, where $L>2$, that have locally Lipschitz gradients. Our contribution can be summarized as follows:
  • In Theorem 5.1 we describe the gradient flow dynamics of deep homogeneous neural networks, with order of homogeneity strictly greater than two, trained with small initializations. Specifically, for square and logistic losses, we show that if the initialization is sufficiently small, then in the early stages of training the weights remain small and either approximately converge in direction to a non-negative KKT point of the constrained NCF defined with respect to the labels in the training set, or are approximately $\mathbf{0}$.
  • For square loss, Theorem 5.3 establishes a similar directional convergence for deep homogeneous neural networks near certain saddle points, provided the weights of neural network have a separable structure. We show that when initialized in a sufficiently small neighborhood of a saddle point, then under gradient flow the weights remain near the saddle point for a sufficiently long time, during which the weights with small norm remain small and either approximately converge in direction along non-negative KKT points of the constrained NCF defined with respect to the residual error at that saddle point, or are approximately $\mathbf{0}$.
To prove our results, we build upon findings from [1] along with crucial new insights; see Section 5.1.1 for details. It is worth noting that we assume neural networks to have locally Lipschitz gradients, which excludes deep ReLU networks. However, it does include deep linear networks [11, 12], and deep neural networks with differentiable and homogeneous activation such as higher powers of ReLU $\phi(x):=\max(x,0)^{p}$ for $p\geq 2$ [13, 14, 15] and polynomials [16, 17]. Moreover, we also discuss the challenges in extending our results for deep ReLU networks, which could be useful in extending our results to ReLU networks.
2 Related Works
Understanding the training dynamics of neural networks have been the subject of numerous investigations [18, 4, 19, 20, 21]. These investigations have revealed two contrasting viewpoints based on the scale of initialization. For large initialization regime, the training dynamics of wide neural networks are effectively captured by a kernel referred to as Neural Tangent Kernel (NTK) [18]. In this regime, the weights remain close to initialization during training, and neural networks behaves like the linearization of neural network around its initialization [4]. On the other hand, in small initialization regime, training dynamics is extremely non-linear, and the change in weights during training is significant [21, 22, 3, 2]. Moreover, studies like [3, 4] have observed improved generalization with decreasing initialization scale for various tasks, which makes understanding the dynamics of neural networks in the small initialization regime crucial for better understanding of practical success of deep networks. While considerable progress has been made on understanding the dynamics of neural networks in the large initialization regime, the theoretical investigations into the small initialization regime has been comparatively limited, primarily due to the extremely non-linear dynamics during training. Such investigations have predominantly focused on linear neural networks [5, 12, 23] and shallow non-linear neural networks [22, 21, 24, 20]. Even for these networks, a comprehensive understanding of the gradient descent dynamics is further limited to diagonal linear networks [5], shallow matrix factorization [25, 26], and two-layer (Leaky) ReLU networks under various simplifying assumptions on the datasets [8, 9, 27, 10, 28]. Our work could be the first step towards a greater understanding of the training dynamics of deep neural networks. The phenomenon of early directional convergence for small initializations was first investigated by [6] for two-layer ReLU neural networks. This was later extended to two-homogeneous neural networks by [1], which includes, in addition to two-layer ReLU neural networks, deep ReLU neural networks with only two trainable layers and single layer squared ReLU neural networks. Other empirical investigations have observed early directional convergence in some deeper networks as well. For example, [29] looks at the early training dynamics of three-layer ReLU neural networks and observe that weights converge along some specific orientations. In [30] authors empirically show a alignment effect in the kernel of the neural network during early stages of the training. However, rigorous investigations thus far have been limited to shallow neural networks. For classification tasks, another line of work studies directional convergence of weights in the later stages of training homogeneous neural networks [31, 32]. In these works, apart from homogeneity, the tail behavior of the loss functions typically used for classification, for example exponential and logistic loss, also plays an important role. Although the problem considered in this paper is different, some theoretical results developed in these works were helpful for us.
3 Preliminaries
We briefly state important notations and describe some relevant preliminary concepts. For any $N\in\mathbb{N}$, we let $[N]=\{1,2,\ldots,N\}$ denote the set of positive integers less than or equal to $N$. We let $\|\cdot\|_{2}$ denote the $\ell_{2}$ norm for a vector and the spectral norm for a matrix. We write $\dot{x}(t)=:\frac{dx(t)}{dt}$, and for the sake of brevity we may remove the independent variable $t$ if it is clear from context. A $k$-dimensional sphere is denoted by $\mathcal{S}^{k-1}$. Suppose $P$ and $Q$ are locally compact metric spaces, such as Euclidean spaces. Then, a function $f:P\rightarrow Q$ is called locally Lipschitz if and only if it is Lipschitz continuous on every compact subset of $Q$. We will assume the neural networks consider in this paper are definable under some o-minimal structure. This is a mild technical assumption and includes deep neural networks with common activation functions such as ReLU, higher powers of ReLUs and polynomials. Such assumptions have mainly been made in the past works, and here as well, to ensure convergence of certain gradient flow trajectories [31].
4 Problem formulation
We adopt a supervised learning framework for training, where we assume $\{\mathbf{x}_{i},y_{i}\}_{i=1}^{n}$ is the training dataset, and let $\mathbf{X}=[\mathbf{x}_{1},\ldots,\mathbf{x}_{n}]\in\mathbb{R}^{d\times n}$ and $\mathbf{y}=[y_{1},\ldots,y_{n}]^{\top}\in\mathbb{R}^{n}$. For a neural network $\mathcal{H}$, its output is denoted by $\mathcal{H}(\mathbf{x};\mathbf{w})$, where $\mathbf{x}$ is the input and $\mathbf{w}\in\mathbb{R}^{k}$ is the vector containing all the weights. We let $\mathcal{H}(\mathbf{X};\mathbf{w})=\left[\mathcal{H}(\mathbf{x}_{1};\mathbf{w}),\ldots,\mathcal{H}(\mathbf{x}_{n};\mathbf{w})\right]^{\top}\in\mathbb{R}^{n}$ be the vector containing the output of neural network for all inputs, and $\mathcal{J}(\mathbf{X};\mathbf{w}):\mathbb{R}^{k}\rightarrow\mathbb{R}^{n}$ denotes the Jacobian of $\mathcal{H}(\mathbf{X};\mathbf{w})$ with respect to $\mathbf{w}$. The following assumptions formalizes the previously mentioned properties of neural networks considered in this paper.
Assumption 1.
We make the following tripartite assumption on the neural network function: (i) For any fixed $\mathbf{x}$, $\mathcal{H}(\mathbf{x};\mathbf{w})$ is locally Lipschitz and is definable under some o-minimal structure that includes polynomials and exponential222 [33] proved the existence of such an o-minimal structure.. (ii) For all $c>0$, $\mathcal{H}(\mathbf{x};c\mathbf{w})=c^{L}\mathcal{H}(\mathbf{x};\mathbf{w}),$ for some $L>2$ (i.e., the network is $L$-homogenous). (iii) The gradient [^3] Throughout the paper, $\nabla\mathcal{H}(\mathbf{x};\mathbf{w})$ denotes the gradient of $\mathcal{H}(\mathbf{x};\mathbf{w})$ with respect to $\mathbf{w}$. $\nabla\mathcal{H}(\mathbf{x};\mathbf{w})$ is locally Lipschitz.
In this paper, we consider the minimization of
$\mathcal{L}\left(\mathbf{w}\right)=\sum_{i=1}^{n}\ell\left(\mathcal{H}(\mathbf{x}_{i};\mathbf{w}),y_{i}\right),$ (1)
where $\ell(\hat{y},y)$ is the loss function and it is either the square loss, $\ell(\hat{y},y)=(\hat{y}-y)^{2}/2,$ or the logistic loss, $\ell(\hat{y},y)=2\ln(1+e^{-\hat{y}y}).$ We minimize (1) using gradient flow, which leads to following differential equation
$\dot{\mathbf{w}}(t)=-\nabla\mathcal{L}(\mathbf{w}(t)),\mathbf{w}(0)=\delta\mathbf{w}_{0},$ (2)
where $\delta$ is a positive scalar that controls the scale of initialization, and $\mathbf{w}_{0}$ is a vector. Following [1], for a fixed vector $\mathbf{z}\in\mathbb{R}^{n}$ and neural network $\mathcal{H}$, the Neural Correlation Function (NCF) is defined here as
$\mathcal{N}_{\mathbf{z},\mathcal{H}}(\mathbf{w})=\mathbf{z}^{\top}\mathcal{H}(\mathbf{X};\mathbf{w}),$
and the constrained optimization problem
$\max_{\|\mathbf{w}\|^{2}_{2}=1}\mathcal{N}_{\mathbf{z},\mathcal{H}}(\mathbf{w})$
is referred to as the constrained NCF.
5 Main Results
5.1 Early Directional Convergence
Our first main result is the following theorem which establishes the approximate directional convergence in the early stages of training $L$-homogeneous neural networks with small initializations, where $L>2$.
Theorem 5.1.
Let $\mathbf{w}_{0}$ be a fixed unit vector. For any $\epsilon\in(0,\eta/2)$, where $\eta$ is a positive constant, there exists $T_{\epsilon},\tilde{B}_{\epsilon}$ and $\overline{\delta}>0$ such that the following holds: for any $\delta\in(0,\overline{\delta})$ and solution $\mathbf{w}(t)$ of (2) for square or logistic loss with initialization $\mathbf{w}(0)=\delta\mathbf{w}_{0}$, we have
$\left\|\mathbf{w}\left({t}\right)\right\|_{2}\leq\tilde{B}_{\epsilon}\delta,\text{ for all }t\in\left[0,\frac{T_{\epsilon}}{\delta^{L-2}}\right].$
Further, for $\overline{T}_{\epsilon}=T_{\epsilon}/\delta^{L-2}$, either
$\|\mathbf{w}(\overline{T}_{\epsilon})\|_{2}\geq\delta\eta/2,\text{ and }\frac{\mathbf{w}(\overline{T}_{\epsilon})^{\top}{\mathbf{u}^{*}}}{\|\mathbf{w}(\overline{T}_{\epsilon})\|_{2}}\geq 1-\left(1+\frac{3}{\eta}\right)\epsilon,$
where ${\mathbf{u}}^{*}$ is a non-negative KKT  [^4] A KKT point $\hat{\mathbf{u}}$ of $\max_{\|\mathbf{u}\|_{2}^{2}=1}\mathcal{N}_{\mathbf{z},\mathcal{H}}(\mathbf{u})$ is referred to as non-negative KKT point if $\mathcal{N}_{\mathbf{z},\mathcal{H}}(\hat{\mathbf{u}})$ is non-negative. point of
$\max_{\|\mathbf{u}\|_{2}^{2}=1}\mathcal{N}_{\mathbf{y},\mathcal{H}}(\mathbf{u})=\mathbf{y}^{\top}\mathcal{H}(\mathbf{X};\mathbf{u}),$ (3)
or
$\|\mathbf{w}(\overline{T}_{\epsilon})\|_{2}\leq 2\delta\epsilon.$
The above theorem first shows that for a given $\epsilon$, we can choose $\delta$ small enough such that for $t\in[0,T_{\epsilon}/\delta^{L-2}]$ the norm of the weights increases at most by a multiplicative factor. Thus, for small $\delta$ the weights will remain small for $t\in[0,T_{\epsilon}/\delta^{L-2}]$. Furthermore, since $L>2$, for smaller $\delta$ the weights will stay small for a longer duration of time. Next, the theorem describes what happens at $\overline{T}_{\epsilon}=T_{\epsilon}/\delta^{L-2}$, the duration for which the norm of the weights remain small. There are two possible scenarios. In the first scenario, the weights approximately converge in direction along a non-negative KKT point of (3), which is the constrained NCF defined with respect to $\mathbf{y}$ and neural network $\mathcal{H}$. Thus, in the early stages of training, the weights of the neural network converge in direction while their norm remains small. In the second scenario, the weights approximately converge to $\mathbf{0}$, which can be observed by noting that $\|\mathbf{w}(\overline{T}_{\epsilon})\|_{2}\leq 2\delta\epsilon$, where we can choose $\epsilon$ to be arbitrarily small. This is in contrast with the first scenario, where $\|\mathbf{w}(\overline{T}_{\epsilon})\|_{2}\geq\delta\eta/2$ and $\eta$ is a constant that does not depend on $\delta$. This shows that in the second scenario, compared to the first scenario, the weights become much smaller. The second scenario arises due to the convergence of the gradient dynamics of the NCF towards $\mathbf{0}$; see the proof for more details. We also note that for a fixed training data and neural network, $\mathbf{w}_{0}$ determines which of the above two scenarios will occur. In addition to this, the positive constant $\eta$ in the above theorem is also determined by $\mathbf{w}_{0}$ and is independent of $\delta$.
5.1.1 Proof Sketch of Theorem 5.1
We next briefly provide proof sketch of the above theorem. We begin by describing the (positive) gradient flow dynamics of the NCF. Recall, for a given vector $\mathbf{z}$ and a neural network $\mathcal{H}$, the NCF is defined as
$\mathcal{N}_{\mathbf{z},\mathcal{H}}(\mathbf{u})=\mathbf{z}^{\top}\mathcal{H}(\mathbf{X};\mathbf{u}).$ (4)
The gradient flow resulting from maximizing the above function is
$\frac{d{\mathbf{u}}}{dt}=\nabla\mathcal{N}_{\mathbf{z},\mathcal{H}}(\mathbf{u}),{\mathbf{u}}(0)=\mathbf{u}_{0},$ (5)
where $\mathbf{u}_{0}$ is the initialization. Note that the NCF may not be bounded from above. Therefore, gradient flow may potentially diverge to infinity. However, the following lemma shows that the gradient flow either converges in direction to a KKT point of the constrained NCF, or converges to $\mathbf{0}$.
Lemma 5.2.
Under 1, for any solution $\mathbf{u}(t)$ of (5), one of the following two possibilities is true.
  1. 1.
    There exists some finite $T^{*}$ such that $\lim_{t\rightarrow T^{*}}\|\mathbf{u}(t)\|_{2}=\infty$. Further, $\lim_{t\rightarrow T^{*}}\frac{\mathbf{u}(t)}{\|\mathbf{u}(t)\|_{2}}$ exists and is equal to a non-negative KKT point of the optimization problem
    $\max_{\|\mathbf{u}\|_{2}^{2}=1}\mathcal{N}_{\mathbf{z},\mathcal{H}}(\mathbf{u})=\mathbf{z}^{\top}\mathcal{H}(\mathbf{X};\mathbf{u}).$ (6)
  2. 2.
    For all $t\geq 0$, $\|\mathbf{u}(t)\|_{2}<\infty,$ and either $\lim_{t\rightarrow\infty}\frac{\mathbf{u}(t)}{\|\mathbf{u}(t)\|_{2}}$ exists or $\lim_{t\rightarrow\infty}{\mathbf{u}(t)}=0.$ If $\lim_{t\rightarrow\infty}\frac{\mathbf{u}(t)}{\|\mathbf{u}(t)\|_{2}}$ exists then it is equal to non-negative KKT point of (6).
To prove directional convergence in the above lemma, we require the neural networks to be definable with respect to an o-minimal structure. Such a requirement allows us to use the unbounded version of Kurdyka-Lojasiewicz inequality proved in [31] for establishing directional convergence. A similar result was demonstrated in [1] for two-homogeneous networks. However, in that instance, the solution $\mathbf{u}(t)$ remains finite for all finite time [^5] See Lemma D.1., whereas in our case, $\mathbf{u}(t)$ may become unbounded at some finite time. We next explain our proof technique for Theorem 5.1 assuming square loss is used for training. The evolution of $\mathbf{w}(t)$ is governed by the differential equation
$\dot{\mathbf{w}}=\sum_{i=1}^{n}\left({y}_{i}-\mathcal{H}(\mathbf{x}_{i};\mathbf{w})\right)\nabla\mathcal{H}(\mathbf{x}_{i};\mathbf{w})=\mathcal{J}(\mathbf{X};\mathbf{w})^{\top}\left(\mathbf{y}-\mathcal{H}(\mathbf{X};\mathbf{w})\right),\mathbf{w}(0)=\delta\mathbf{w}_{0},$ (7)
where (as mentioned) $\delta>0$, and recall $\mathcal{J}(\mathbf{X};\mathbf{w}):\mathbb{R}^{k}\rightarrow\mathbb{R}^{n}$ is the Jacobian of the function $\mathcal{H}(\mathbf{X};\mathbf{w})$ with respect to $\mathbf{w}$. We define $\mathbf{s}(t)=\frac{1}{\delta}\mathbf{w}\left(\frac{t}{\delta^{L-2}}\right)$; then, $\mathbf{s}(0)=\mathbf{w}_{0}$. Further, by using the $L$-homogeneity of $\mathcal{H}\left(\mathbf{X};\mathbf{w}\right)$ and $(L-1)$-homogeneity of $\mathcal{J}(\mathbf{X};\mathbf{w})$ (Lemma B.1), we can derive the evolution of $\mathbf{s}(t)$ as follows
$\frac{d\mathbf{s}}{dt}=\frac{d}{dt}\left(\frac{1}{\delta}\mathbf{w}\left(\frac{t}{\delta^{L-2}}\right)\right)$ $=\frac{1}{\delta^{L-1}}\dot{\mathbf{w}}\left(\frac{t}{\delta^{L-2}}\right)$
$=\frac{1}{\delta^{L-1}}\mathcal{J}\left(\mathbf{X};\mathbf{w}\left(\frac{t}{\delta^{L-2}}\right)\right)^{\top}\left(\mathbf{y}-\mathcal{H}\left(\mathbf{X};\mathbf{w}\left(\frac{t}{\delta^{L-2}}\right)\right)\right)$
$=\mathcal{J}\left(\mathbf{X};\frac{1}{\delta}\mathbf{w}\left(\frac{t}{\delta^{L-2}}\right)\right)^{\top}\left(\mathbf{y}-\delta^{L}\mathcal{H}\left(\mathbf{X};\frac{1}{\delta}\mathbf{w}\left(\frac{t}{\delta^{L-2}}\right)\right)\right).$
Overall, this implies
$\frac{d\mathbf{s}}{dt}$ $=\mathcal{J}\left(\mathbf{X};\mathbf{s}\right)^{\top}(\mathbf{y}-\delta^{L}\mathcal{H}\left(\mathbf{X};\mathbf{s}\right)),\mathbf{s}(0)=\mathbf{w}_{0}.$ (8)
Now, we compare the dynamics for $\mathbf{s}(t)$ in (8) with the dynamics of gradient flow of the NCF defined with respect to $\mathbf{y}$ and neural network $\mathcal{H}$,
$\dot{\mathbf{u}}=\nabla\mathcal{N}_{\mathbf{z},\mathcal{H}}(\mathbf{u})=\mathcal{J}(\mathbf{X};\mathbf{u})^{\top}\mathbf{y},{\mathbf{u}}(0)=\mathbf{w}_{0}.$ (9)
Observe that if $\delta=0$, then the two dynamics are exactly the same, and since $\mathbf{s}(0)=\mathbf{w}_{0}=\mathbf{u}(0)$, we have that $\mathbf{s}(t)$ and $\mathbf{u}(t)$ will also be same. For $\delta>0$, there is an extra term $\delta^{L}\mathcal{H}\left(\mathbf{X};\mathbf{s}\right)$ present in (8). For small $\delta$, we show that this extra term is small and while $\mathbf{s}(t)$ and $\mathbf{u}(t)$ are not exactly the same, the difference between them is small, i.e., $\|\mathbf{s}(t)-\mathbf{u}(t)\|_{2}$ is small. Moreover, we show that the difference can be made sufficiently small for a sufficiently long time by choosing $\delta$ small enough. We combine this fact with Lemma 5.2 to prove Theorem 5.1. Specifically, from Lemma 5.2, we know that $\mathbf{u}(t)$ will either converge to $\mathbf{0}$ or converge in direction to a KKT point of the constrained NCF. Therefore, we first choose a time (say $T$) such that $\mathbf{u}(T)$ has approximately converged in direction to a KKT point of the constrained NCF or is in the neighborhood of $\mathbf{0}$. Then, we choose $\delta$ sufficiently small such that $\|\mathbf{s}(t)-\mathbf{u}(t)\|_{2}$ is small for all $t\in[0,T]$. Since $\mathbf{s}(t)=\frac{1}{\delta}\mathbf{w}\left(\frac{t}{\delta^{L-2}}\right)$, we have that $\|\frac{1}{\delta}\mathbf{w}\left(\frac{T}{\delta^{L-2}}\right)-\mathbf{u}(T)\|_{2}$ is also small. Finally, since $\delta$ is a positive scalar, and $\mathbf{u}(T)$ approximately converges in direction to a KKT point of the constrained NCF or is approximately $\mathbf{0}$, it follows that $\mathbf{w}\left(\frac{T}{\delta^{L-2}}\right)$ also approximately converges in direction to a KKT point of the constrained NCF or is approximately $\mathbf{0}$. As a quick aside, we note that there are two key aspects in the definition $\mathbf{s}(t)=\frac{1}{\delta}\mathbf{w}\left(\frac{t}{\delta^{L-2}}\right)$. First, we have divided by $\delta$. We know that $\delta$ controls the scale of initialization and is assumed to be small. Dividing by $\delta$ removes that scaling and makes $\mathbf{s}(t)$ a constant scale vector. More importantly, dividing by $\delta$ does not alter the direction, i.e., the direction of $\mathbf{s}(t)$ and $\mathbf{w}(t)$ are always the same. Thus, establishing directional convergence of $\mathbf{s}(t)$ will also imply directional convergence of $\mathbf{w}(t)$. Second, we rescale the time by $\delta^{L-2}$. To understand the importance of this time rescaling, let us look at the magnitude of the gradient of the loss at initialization in terms of $\delta$. Since $\delta$ is small and $\mathbf{w}_{0}$ has unit norm, $\|\mathcal{J}(\mathbf{X};\mathbf{w}(0))^{\top}\left(\mathbf{y}-\mathcal{H}(\mathbf{X};\mathbf{w}(0))\right)\|_{2}\approx\|\mathcal{J}(\mathbf{X};\mathbf{w}(0))^{\top}\mathbf{y}\|_{2}=O(\delta^{L-1})$, where in the last equality we used $(L-1)$-homogeneity of the Jacobian. We observe that for initialization of scale $\delta$, the gradient scales as $\delta^{L-1}$. Thus, in the early stages the gradient will have negligible effect on the weights, in terms of both magnitude and direction. Moreover, based on the relative scaling of the gradient and the initialization, the gradient will only start impacting the weights after $O(1/\delta^{L-2})$ time has elapsed. Hence, for smaller initialization, the gradient flow will take longer to make an impact on the weights. Thus, in a way, rescaling time by $\delta^{L-2}$ brings the time of $\mathbf{s}(t)$ to a “constant scale”. We also highlight the importance of homogeneity. It allows the above two changes to interact nicely with the gradient flow and results in dynamics of $\mathbf{s}(t)$ that are essentially independent of $\delta$, provided $\delta$ is small. Finally, in comparison to [1], which studies early directional convergence for two-homogeneous neural networks, the main conceptual innovation in this paper is the time rescaling by $\delta^{L-2}$. In fact, the proof in that work relies on showing that the difference between $\mathbf{w}(t)/\delta$ and $\mathbf{u}(t)$ remains small for a sufficiently long time. To see a motivation for this, note that we can get (8) from (7) for $L=2$ as well. However, for $L=2$, the time rescaling factor $\delta^{L-2}$ will become unity, yielding $\mathbf{s}(t)=\mathbf{w}(t)/\delta$. Now, since the gradient flow dynamics for the NCF of two-homogeneous neural networks also converges to a KKT point of the constrained NCF, to establish early directional convergence one must show that the difference between $\mathbf{w}(t)/\delta$ and $\mathbf{u}(t)$ remains small for a sufficiently long time. Another notable difference with [1] is in the gradient flow dynamics of the NCF. As mentioned previously, from Lemma 5.2, we know that for $L>2$, the gradient flow dynamics of the NCF can become unbounded at some finite time, whereas for $L=2$, the dynamics remains finite for finite time. This difference necessitates a more careful analysis of the early stages of training in order to ensure directional convergence for deep homogeneous neural networks.
5.1.2 A Closer Look at the Weights
In Theorem 5.1 we established that in the early stages of training, for sufficiently small initialization, the weights of neural network approximately converge in direction along the KKT points of the NCF while their norm remained small. In this section, we use a toy example to highlight the emergence of additional structural properties in the weights during the early stage of training that are not explained by Theorem 5.1. We train the following $4-$homogeneous neural network $\mathcal{H}(\mathbf{x};\mathbf{W}_{1},\mathbf{W}_{2})=\mathbf{1}^{\top}\max(\mathbf{W}_{2}\mathbf{W}_{1}\mathbf{x},\mathbf{0})^{2}$, where $\mathbf{W}_{1}\in\mathbb{R}^{20\times 10}$, $\mathbf{W}_{2}\in\mathbb{R}^{20\times 20}$ and $\mathbf{1}\in\mathbb{R}^{20}$ is the all-one vector. Here, we train to minimize the square loss with respect to the output of a smaller neural network *$\mathcal{H}^{*}(\mathbf{x};\mathbf{W}_{1}^{*},\mathbf{W}_{2}^{*})=\mathbf{1}^{\top}\max(\mathbf{W}_{2}^{*}\mathbf{W}_{1}^{*}\mathbf{x},\mathbf{0})^{2}$, where the entries of $\mathbf{W}_{1}^{*}\in\mathbb{R}^{2\times 10}$, $\mathbf{W}_{2}^{*}\in\mathbb{R}^{3\times 2}$ are drawn from the standard normal distribution and $\mathbf{1}\in\mathbb{R}^{3}$ is the all-one vector. The training data has 100 points sampled uniformly from unit sphere in $\mathbb{R}^{10}$. Let $\mathbf{w}$ be a vector containing all the entries of $\mathbf{W}_{1},\mathbf{W}_{2}$, and $\tilde{\mathbf{w}}=\mathbf{w}/\|\mathbf{w}\|_{2}.$ For training, we use gradient descent with step-size $5\cdot 10^{-3}$, and the initialization $\mathbf{w}(0)=\delta\mathbf{w}_{0}$, where $\delta=0.1$ and $\mathbf{w}_{0}$ is a random unit norm vector. We train for 99100 iterations, and in Figure 0(a), plot the evolution of training loss and $\ell_{2}-$norm of the weights. It is evident from the figure that loss almost remains the same, and the norm of the weights has increased but only by a constant factor and is still small. It can be easily shown (see Lemma B.3) that a unit norm vector is a KKT point of the constrained NCF if the vector and the gradient of the NCF at that vector are parallel. Therefore, to measure the directional convergence of $\mathbf{w}$ to the KKT point of the constrained NCF, in Figure 0(b) we plot the evolution of the inner product between $\tilde{\mathbf{w}}(t)$ and $\nabla\mathcal{N}(\tilde{\mathbf{w}}(t))/\|\nabla\mathcal{N}(\tilde{\mathbf{w}}(t))\|_{2}$, where $\mathcal{N}(\cdot)$ denotes the NCF for this example. Clearly, in accordance with Theorem 5.1, the weights have approximately converged in direction to a KKT point of the constrained NCF.
Figure 1: Early training dynamics of a squared-ReLU neural network. Panel $(a)$: the evolution of training loss (normalized with respect to loss at initialization) and the $\ell_{2}$-norm of all the weights with iterations. Panel $(b)$: the evolution of $\nabla\mathcal{N}(\tilde{\mathbf{w}}(t))^{\top}\tilde{\mathbf{w}}(t)/\|\nabla\mathcal{N}(\tilde{\mathbf{w}}(t))\|_{2}$ (a measure of directional convergence of $\mathbf{w}(t)$). We observe that the loss has barely changed, and the norm of weights has increased by a constant factor but remains small. Also, the weights approximately converge in direction to a KKT point of the constrained NCF.
We next take a closer look at the weights of the neural networks at iteration 99100, i.e., when we stop training and the weights have approximately converged in direction to a KKT point of the constrained NCF. In Figure 2, we plot each row of $\mathbf{W}_{1}/\|\mathbf{w}\|_{2}$ and $\mathbf{W}_{2}/\|\mathbf{w}\|_{2}$ at initialization and at iteration 99100. At initialization the rows of $\mathbf{W}_{1}/\|\mathbf{w}\|_{2}$ and $\mathbf{W}_{2}/\|\mathbf{w}\|_{2}$ appear to be randomly oriented, as expected. At iteration 99100, all the rows seem to be oriented along certain specific directions and there is a low-rank structure among $\mathbf{W}_{1}/\|\mathbf{w}\|_{2}$ and $\mathbf{W}_{2}/\|\mathbf{w}\|_{2}$. The low-rank structure can be further verified by observing the singular values of $\mathbf{W}_{1}/\|\mathbf{w}\|_{2}$ and $\mathbf{W}_{2}/\|\mathbf{w}\|_{2}$, depicted in Figure 3.
Figure 2: Each curve in the plots correspond to a row of $\mathbf{W}_{1}/\|\mathbf{w}\|_{2}$ or $\mathbf{W}_{2}/\|\mathbf{w}\|_{2}$ at different stages of training. Panel $(a)$: $\mathbf{W}_{1}/\|\mathbf{w}\|_{2}$ at initialization. Panel $(b)$: $\mathbf{W}_{1}/\|\mathbf{w}\|_{2}$ at iteration 99100. Panel $(c)$: $\mathbf{W}_{2}/\|\mathbf{w}\|_{2}$ at initialization. Panel $(d)$: $\mathbf{W}_{2}/\|\mathbf{w}\|_{2}$ at iteration 99100.
Figure 3: Panel $(a)$: singular values of $\mathbf{W}_{1}/\|\mathbf{w}\|_{2}$ at iteration 99100. Panel $(b)$: singular values of $\mathbf{W}_{2}/\|\mathbf{w}\|_{2}$ at iteration 99100.
We clarify that the results of Theorem 5.1 neither explain how such low-rank structures emerge in the weights, nor characterizes such low-rank structures. We leave it as an important future research direction to explain the emergence of these low-rank structures. Notably, similar low-rank behavior under small initialization was observed by [29] for three-layer ReLU networks (which is 3-homogeneous). Also, [30] has empirically shown emergence of low-rank structures in the kernel learned by neural networks during the early phase of training.
5.1.3 Challenges for Non-differentiable Neural Networks
Recall that Theorem 5.1 holds for neural networks with locally Lipschitz gradient, and thus excludes ReLU neural networks. We first provide an empirical evidence that early directional convergence also occurs in ReLU networks, and then briefly describe the challenges in extending Theorem 5.1 to ReLU networks. Let $\sigma(x)=\max(x,0)$. We train the following $3-$homogeneous ReLU network $\mathcal{H}(\mathbf{x};\mathbf{W}_{1},\mathbf{W}_{2},\mathbf{W}_{3})=\mathbf{1}^{\top}\sigma(\mathbf{W}_{3}\sigma(\mathbf{W}_{2}\sigma(\mathbf{W}_{1}\mathbf{x})))$, where $\mathbf{W}_{1}\in\mathbb{R}^{20\times 10}$, $\mathbf{W}_{2},\mathbf{W}_{3}\in\mathbb{R}^{20\times 20}$ and $\mathbf{1}\in\mathbb{R}^{20}$ is the all-one vector. For training, we minimize the square loss with respect to the output of a smaller neural network *$\mathcal{H}^{*}(\mathbf{x};\mathbf{W}_{1}^{*},\mathbf{W}_{2}^{*},\mathbf{W}_{3}^{*})=\mathbf{1}^{\top}\sigma(\mathbf{W}_{3}^{*}\sigma(\mathbf{W}_{2}^{*}\sigma(\mathbf{W}_{1}^{*}\mathbf{x})))$, where the entries of of $\mathbf{W}_{1}^{*}\in\mathbb{R}^{2\times 10}$, $\mathbf{W}_{2}^{*}\in\mathbb{R}^{3\times 2},\mathbf{W}_{3}^{*}\in\mathbb{R}^{3\times 3}$ are drawn from standard normal distribution and $\mathbf{1}\in\mathbb{R}^{3}$ is the all-one vector. The training data has 100 points sampled uniformly from unit sphere in $\mathbb{R}^{10}$. We define $\mathbf{w}$ and $\tilde{\mathbf{w}}$ in a similar way as in the previous experiment, and $\mathcal{N}(\mathbf{w})$ denotes the NCF for this case. The network is trained for 80000 iterations using gradient descent with step-size $10^{-3}$, and the initialization $\mathbf{w}(0)=\delta\mathbf{w}_{0}$, where $\delta=0.01$ and $\mathbf{w}_{0}$ is a random unit norm vector. From Figure 3(a), we observe that the training loss does not change much, and the norm of the weights increases by a multiplicative factor but remains small. Also, from Figure 3(b), it is clear that the weights approximately converge in direction to a KKT point of the constrained NCF. Furthermore, similar to the previous experiment, low-rank structure emerges among the weights of neural networks; see Figure 5.
Figure 4: Early training dynamics of a ReLU neural network. Panel $(a)$: the evolution of training loss (normalized with respect to loss at initialization) and the $\ell_{2}$-norm of all the weights with iterations. Panel $(b)$: the evolution of $\partial\mathcal{N}(\tilde{\mathbf{w}}(t))^{\top}\tilde{\mathbf{w}}(t)/\|\partial\mathcal{N}(\tilde{\mathbf{w}}(t))\|_{2}$ (a measure of directional convergence of $\mathbf{w}(t)$). The loss does not change much, and the norm of weights has increased by a constant factor but remains small. Also, the weights approximately converge in direction to a KKT point of the constrained NCF.
We next describe the difficulty in extending Theorem 5.1 to ReLU networks. Consider a neural network that satisfies 1, except that the gradient is not assumed to be locally Lipschitz (this will include ReLU networks). Then, assuming square loss, we can define the gradient flow dynamics using the Clarke subdifferential as follows
$\dot{\mathbf{w}}\in\sum_{i=1}^{n}\left({y}_{i}-\mathcal{H}(\mathbf{x}_{i};\mathbf{w})\right)\partial\mathcal{H}(\mathbf{x}_{i};\mathbf{w}),\mathbf{w}(0)=\delta\mathbf{w}_{0}.$ (10)
It is also well-known that the Clarke differential of an $L$-homogeneous function is $(L-1)$-homogeneous [32, Theorem B.2]. Thus, if we define $\mathbf{s}(t)=\frac{1}{\delta}\mathbf{w}\left(\frac{t}{\delta^{L-2}}\right)$, then similar to (8), we can show that $\mathbf{s}(t)$ will evolve according to
$\dot{\mathbf{s}}$ $\in\sum_{i=1}^{n}({y}_{i}-\delta^{L}\mathcal{H}\left(\mathbf{x}_{i};\mathbf{s}\right))\partial\mathcal{H}\left(\mathbf{x}_{i};\mathbf{s}\right),{\mathbf{s}}(0)=\mathbf{w}_{0}.$ (11)
Further, the gradient flow for the NCF can be defined by the differential inclusion
$\dot{\mathbf{u}}\in\sum_{i=1}^{n}y_{i}\partial\mathcal{H}(\mathbf{x}_{i};\mathbf{u}),{\mathbf{u}}(0)=\mathbf{w}_{0}.$ (12)
Following the proof sketch of Theorem 5.1 described in Section 5.1.1, to show approximate directional converegence of $\mathbf{w}(t)$, we would need to first show that for small enough $\delta$, the solutions of (12) and (11) will be sufficiently small for sufficiently long time. From comparing (12) and (11), it is clear that for small $\delta$, the two solutions will be close in the beginning. However, we can not claim that the solutions will remain close for sufficiently long time by choosing small enough $\delta$. In fact, the proof of the analogous result in the context of Theorem 5.1 crucially relied on gradients being locally Lipschitz, which is not the case here. We believe that proving this would be a crucial step towards establishing directional convergence for ReLU networks, and defer it for future investigations.
Figure 5: Panel $(a)$: $\mathbf{W}_{1}/\|\mathbf{w}\|_{2}$ at iteration 80000. Panel $(b)$: $\mathbf{W}_{2}/\|\mathbf{w}\|_{2}$ at iteration 80000. Panel $(c)$: $\mathbf{W}_{3}/\|\mathbf{w}\|_{2}$ at iteration 80000. Each curve in the plots correspond to a row of the corresponding weights matrix.
5.2 Directional Convergence Near Saddle Points
For small initializations and square loss, several recent studies have also observed a so-called saddle-to-saddle dynamics during training of neural networks; see, e.g., [34, 35], where it is observed that during training the weights of neural networks passes through a sequence of saddle points. Moreover, recent theoretical results have further revealed that only a certain number of weights are non-zero at saddle point encountered during training [8, 35]. Motivated by these observations, in this section we analyze the gradient flow dynamics near certain saddle points, specifically those saddle points where only a certain number of weights are non-zero while the others are zero. It is worth noting that [1] also analyzed the gradient flow dynamics near similar types of saddle points for two-homogeneous neural networks. In this section, we assume that the neural network has a separable structure. Specifically, we assume that the weights of the neural network $\mathbf{w}$ can be partitioned into two sets, that is, $\mathbf{w}=\left[\mathbf{w}_{n},\mathbf{w}_{z}\right]$, such that the output of the neural network can be written as
$\mathcal{H}(\mathbf{x};\mathbf{w})=\mathcal{H}_{n}(\mathbf{x};\mathbf{w}_{n})+\mathcal{H}_{z}(\mathbf{x};\mathbf{w}_{z}),$
where $\mathcal{H}_{n}(\mathbf{x};\mathbf{w}_{n})$ and $\mathcal{H}_{z}(\mathbf{x};\mathbf{w}_{z})$ are assumed to satisfy 1. Now, let $\mathcal{H}_{n}(\mathbf{X};\mathbf{w}_{n})$ and $\mathcal{H}_{z}(\mathbf{X};\mathbf{w}_{z})$ be defined similarly to $\mathcal{H}(\mathbf{X};\mathbf{w})$, then for square loss, the training loss is
$\mathcal{L}\left(\mathbf{w}_{n},\mathbf{w}_{z}\right)=\frac{1}{2}\left\|\mathcal{H}_{n}(\mathbf{X};\mathbf{w}_{n})+\mathcal{H}_{z}(\mathbf{X};\mathbf{w}_{z})-\mathbf{y}\right\|^{2}.$ (13)
We make the following assumption on the saddle point of (13) considered in this section.
Assumption 2.
$\{\overline{\mathbf{w}}_{n},\overline{\mathbf{w}}_{z}\}$ is a saddle point of (13) such that
$\|\overline{\mathbf{w}}_{n}\|_{2}\in[m,M],\text{ and }\|\overline{\mathbf{w}}_{z}\|_{2}=0,$ (14)
where $m<M$, and $m,M$ are positive constants. Furthermore, there exists $\gamma>0$ such that
$\left\langle\mathbf{w}_{n}-\overline{\mathbf{w}}_{n},\nabla_{\mathbf{w}_{n}}\mathcal{L}(\mathbf{w}_{n},\mathbf{0})\right\rangle\geq 0,\text{ if }\|\mathbf{w}_{n}-\overline{\mathbf{w}}_{n}\|_{2}\leq\gamma.$ (15)
The partition of weights and the above assumption is particularly motivated by aforementioned observation that at saddle points encountered during training, only a certain number of weights are non-zero while others are zero. Clearly, $\overline{\mathbf{w}}_{n}(\overline{\mathbf{w}}_{z})$ represents the set of weights that are non-zero (zero). Additionally, the inequality in (15) is a technical condition to ensure that the gradient flow remains near the saddle point for a sufficiently long time if initialized near the saddle point. We will explain the motivation behind this condition after discussing our main result of this section. The following theorem describes the gradient flow dynamics of (13) when initialized in the neighborhood of a saddle point.
Theorem 5.3.
Let $\{\overline{\mathbf{w}}_{n},\overline{\mathbf{w}}_{z}\}$ satisfy 2 and let $\overline{\mathbf{y}}={\mathbf{y}}-\mathcal{H}_{n}(\mathbf{X};\overline{\mathbf{w}}_{n}).$ Suppose $\mathbf{p}_{n},\mathbf{p}_{z}$ are fixed unit norm vectors. For any $\epsilon\in(0,\eta/2)$, where $\eta$ is a positive constant, there exist $T_{\epsilon},\tilde{B}_{\epsilon}$, $\tilde{C}_{\epsilon}$ and $\overline{\delta}>0$ such that the following holds: for any $\delta\in(0,\overline{\delta})$ and gradient flow solution $\{\mathbf{w}_{n}(t),\mathbf{w}_{z}(t)\}$ of (13) initialized at $\{\overline{\mathbf{w}}_{n}+\delta\mathbf{p}_{n},\overline{\mathbf{w}}_{z}+\delta\mathbf{p}_{z}\}$, we have
$\|\mathbf{w}_{n}(t)-\overline{\mathbf{w}}_{n}\|_{2}\leq\tilde{C}_{\epsilon}\delta,\text{ and }\|\mathbf{w}_{z}(t)-\overline{\mathbf{w}}_{z}\|_{2}\leq\tilde{B}_{\epsilon}\delta,\text{ for all }t\in\left[0,\frac{T_{\epsilon}}{\delta^{L-2}}\right].$ (16)
Further, for $\overline{T}_{\epsilon}=T_{\epsilon}/\delta^{L-2}$, either
$\|\mathbf{w}_{z}(\overline{T}_{\epsilon})\|_{2}\geq\delta\eta/2,\text{ and }\frac{\mathbf{w}_{z}(\overline{T}_{\epsilon})^{\top}{\mathbf{u}}^{*}}{\|\mathbf{w}_{z}(\overline{T}_{\epsilon})\|_{2}}\geq 1-\left(1+\frac{3}{\eta}\right)\epsilon,$
where ${\mathbf{u}}^{*}$ is a non-negative KKT point of
$\max_{\|\mathbf{u}\|_{2}^{2}=1}\mathcal{N}_{\bar{\mathbf{y}},\mathcal{H}_{z}}(\mathbf{u})=\overline{\mathbf{y}}^{\top}\mathcal{H}_{z}(\mathbf{X};\mathbf{u}),$ (17)
or
$\|\mathbf{w}_{z}(\overline{T}_{\epsilon})\|_{2}\leq 2\delta\epsilon.$
In the above theorem, we choose initialization to be near a saddle point of (13) that satisfies 2, where $\delta>0$ controls the distance of initialization from the saddle point. Since $\overline{\mathbf{w}}_{z}=\mathbf{0}$ and $\mathcal{H}_{z}$ is homogeneous, we have $\mathcal{H}_{z}(\mathbf{x};\overline{\mathbf{w}}_{z})=\mathbf{0}$, Therefore, the vector $\overline{\mathbf{y}}$ represents the residual error at the saddle point $\{\overline{\mathbf{w}}_{n},\overline{\mathbf{w}}_{z}\}$, and plays the same role as $\mathbf{y}$ in Theorem 5.1. In the first part we show that for a given $\epsilon$, we can choose $\delta$ sufficiently small such that gradient flow remains near the saddle point for $t\leq T_{\epsilon}/\delta^{L-2}$. We next describe what happens at $\overline{T}_{\epsilon}=T_{\epsilon}/\delta^{L-2}$. Similar to Theorem 5.1, the weights of small magnitude $\mathbf{w}_{z}$ remains small in norm and either approximately converge in direction to a KKT point of (17), which is the constrained NCF defined with respect to $\overline{\mathbf{y}}$ and $\mathcal{H}_{z}$, or are approximately $\mathbf{0}$. Furthermore, for a fixed training data and neural network, $\mathbf{p}_{z}$ determines which of the above two scenarios will occur, and the positive constant $\eta$ in the above theorem is also determined by $\mathbf{p}_{z}$ and is independent of $\delta$. The proof of the above theorem is similar to proof of Theorem 5.1; see the Appendix C for details. We now turn towards explaining the motivation for (15) in 2. An important step towards establishing directional convergence of $\mathbf{w}_{z}$ is to show that $\mathbf{w}_{n}(t)$ remains near $\overline{\mathbf{w}}_{n}$ and $\mathbf{w}_{z}(t)$ remains small in norm for a sufficiently long time, i.e., (16) holds. Notably, the desired amount of time is inversely proportional to $\delta^{L-2}$. Thus, the closer the initialization is to the saddle point, the longer (16) needs to hold for. Let us assume $\mathbf{w}_{z}(t)$ remains small in norm. In that case, $\mathbf{w}_{n}(t)$ approximately evolves under the gradient flow of $\mathcal{L}(\mathbf{w}_{n},\mathbf{0})$. We use the result of the following lemma to describe the gradient flow dynamics of $\mathcal{L}(\mathbf{w}_{n},\mathbf{0})$ near $\overline{\mathbf{w}}_{n}$.
Lemma 5.4.
Let $\{\overline{\mathbf{w}}_{n},\overline{\mathbf{w}}_{z}\}$ be a saddle point of (13) such that $\|\overline{\mathbf{w}}_{n}\|_{2}\in[m,M],\text{ and }\|\overline{\mathbf{w}}_{z}\|_{2}=0,$ where $m,M$ are positive constants as above. Then, $\overline{\mathbf{w}}_{n}$ is a saddle point of
$\mathcal{L}\left(\mathbf{w}_{n},\mathbf{0}\right)=\frac{1}{2}\left\|\mathcal{H}_{n}(\mathbf{X};\mathbf{w}_{n})-\mathbf{y}\right\|^{2}.$ (18)
If $\overline{\mathbf{w}}_{n}$ satisfies (15) for some $\gamma>0$, then $\overline{\mathbf{w}}_{n}$ is a local minimum of (18). Further, if $\mathcal{H}_{n}(.)$ is twice continuously differentiable, and $\overline{\mathbf{w}}_{n}$ is a local minimum of (18) such that $\nabla_{\mathbf{w}_{n}}^{2}\mathcal{L}\left(\overline{\mathbf{w}}_{n},\mathbf{0}\right)\succ\mathbf{0}$, then $\overline{\mathbf{w}}_{n}$ satisfies (15) for some $\gamma>0$.
From the above lemma, we have that $\overline{\mathbf{w}}_{n}$ is a saddle point of $\mathcal{L}(\mathbf{w}_{n},\mathbf{0})$. However, it may not be stable. That is, if $\mathbf{w}_{n}(t)$ is initialized near $\overline{\mathbf{w}}_{n}$, then $\mathbf{w}_{n}(t)$ may not remain near $\overline{\mathbf{w}}_{n}$ for the desired amount of time. However, under (15) $\overline{\mathbf{w}}_{n}$ is a local minimum $\mathcal{L}(\mathbf{w}_{n},\mathbf{0})$, and thus enables stable behavior of gradient flow near $\overline{\mathbf{w}}_{n}$. Hence, (15) ensures that $\mathbf{w}_{n}(t)$ remains near $\overline{\mathbf{w}}_{n}$ for the desired amount of time. We conclude this section by highlighting a major drawback in Theorem 5.3. We have assumed that the weights of the neural network $\mathbf{w}$ can be partitioned into two sets such that the output of the neural network satisfies $\mathcal{H}(\mathbf{x};\mathbf{w})=\mathcal{H}_{n}(\mathbf{x};\mathbf{w}_{n})+\mathcal{H}_{z}(\mathbf{x};\mathbf{w}_{z})$. While this assumption is crucial for our analysis, it excludes fully connected deep neural networks, since such a partition is not possible for those networks. Therefore, Theorem 5.3 is limited to neural networks that have a separable structure. A main reason preventing us from studying gradient flow dynamics near saddle points of fully connected deep neural networks is the lack of theoretical results for such cases. So far, saddle-to-saddle dynamics has only been rigorously studied for shallow networks. Therefore, we do not have a precise knowledge about the saddle points encountered when training fully connected networks. We believe some kind of structural assumptions on the saddle points would be necessary to extend Theorem 5.3 to fully connected networks. We defer investigations along these lines to a future work.
6 Conclusion and Future Directions
In this paper, we studied the gradient flow dynamics resulting from training deep homogeneous neural networks with small initializations, and established the approximate directional convergence of the weights in the early stages of training. For square loss, we also established similar directional convergence near certain saddle points. However, our results here are not applicable for ReLU networks, so one important future direction would be to rigorously establish similar directional convergence for ReLU networks. Using simple experiments, we highlight the directional convergence of weights, as well as a low-rank structure in the weights of neural network during early stages of training. This latter phenomenon is not explained by our work, and is an intriguing area for future exploration. Another interesting future direction would be to study the entire gradient dynamics of deep neural networks, and our work here could potentially aid in such an endeavor. A crucial step for this would be to extend our results at saddle points for more general fully-connected neural networks.
Appendix A Key Lemmata
In this section, we state some key lemmata that are important for our results.
A.1 Kurdyka-Lojasiewicz Inequality for Definable Functions
The Kurdyka-Lojasiewicz Inequality is useful for demonstrating convergence of bounded gradient flow trajectories by establishing the existence of a desingularizing function, which is formally defined as follows.
Definition A.1.
A function $\Psi:[0,\nu)\rightarrow\mathbb{R}$ is called a desingularizing function when $\Psi$ is continuous on $[0,\nu)$ with $\Psi(0)=0$, and it is continuously differentiable on $(0,\nu)$ with $\Psi^{\prime}>0.$
In establishing the directional convergence of gradient flow trajectories of the neural correlation function, where the trajectories may not be bounded, we will utilize the following unbounded version of the Kurdyka-Lojasiewicz inequality.
Lemma A.1.
[31, Lemma 3.6] Let $f$ be a locally Lipschitz function, definable under an o-minimal structure, and having an open domain $\mathbf{D}\subset\{\mathbf{x}|\|\mathbf{x}\|_{2}>1\}$. For any $c,\eta>0$, there exists a $\nu>0$ and a definable desingularizing function $\Psi$ on $[0,\nu)$ such that
$\Psi^{\prime}(f(\mathbf{x}))\|\mathbf{x}\|_{2}\|{\nabla}f(\mathbf{x})\|_{2}\geq 1,\text{ if $f(\mathbf{x})\in(0,\nu)$ and }\|{\nabla}_{\perp}f(\mathbf{x})\|_{2}\geq c\|\mathbf{x}\|_{2}^{\eta}\|{\nabla}_{r}f(\mathbf{x})\|_{2},$
where ${\nabla}_{r}f(\mathbf{x})=\left\langle{\nabla}f(\mathbf{x}),\frac{\mathbf{x}}{\|\mathbf{x}\|_{2}}\right\rangle\frac{\mathbf{x}}{\|\mathbf{x}\|_{2}}$ and ${\nabla}_{\perp}f(\mathbf{x})={\nabla}f(\mathbf{x})-{\nabla}_{r}f(\mathbf{x})$.
A.2 Euler’s Theorem for Homogeneous Functions
The next lemma states two important properties of homogeneous functions.
Lemma A.2.
([32, Theorem B.2], [31, Lemma C.1]) Let $F:\mathbb{R}^{k}\rightarrow\mathbb{R}$ be locally Lipschitz, differentiable, and $L-$positively homogeneous for some $L>0$. Then,
  1. 1.
    For any $\mathbf{w}\in\mathbb{R}^{k}$ and $c\geq 0$,
    $\nabla F(c\mathbf{w})=c^{L-1}\nabla F(\mathbf{w}).$
  2. 2.
    For any $\mathbf{w}\in\mathbb{R}^{k}$,
    $\mathbf{w}^{\top}\nabla F(\mathbf{w})=LF(\mathbf{w}).$
A.3 Gronwall’s Inequality
In our proofs, we will frequently use Gronwall’s inequality.
Lemma A.3.
Let $\alpha,\beta,u$ be real-valued functions defined on an interval $[a,b]$, where $\beta$ and $u$ are continuous and $\min(\alpha,0)$ is integrable on every closed and bounded sub-interval of $[a,b]$.
  • If $\beta$ is non-negative and if $u$ satisfies the integral inequality
[…]
    then
[…]
  • If, in addition, the function $\alpha$ is non-decreasing, then
[…]
We will mainly be using the second part of the above lemma.
Appendix B Proofs Omitted from Section 5.1
B.1 Proof of Lemma 5.2
Recall, for a given vector $\mathbf{z}$ and a neural network $\mathcal{H}$, the NCF is defined as
$\mathcal{N}_{\mathbf{z},\mathcal{H}}(\mathbf{u})=\mathbf{z}^{\top}\mathcal{H}(\mathbf{X};\mathbf{u}).$ (19)
In this section, for the sake of conciseness, we will use $\mathcal{N}(\mathbf{u})$ in place of $\mathcal{N}_{\mathbf{z},\mathcal{H}}(\mathbf{u})$. Further, $\mathbf{u}(t)$ satisfies for all $t\geq 0$
$\frac{d{\mathbf{u}}}{dt}=\nabla\mathcal{N}(\mathbf{u})=\mathcal{J}(\mathbf{X};\mathbf{u})^{\top}\mathbf{z},{\mathbf{u}}(0)=\mathbf{u}_{0},$ (20)
where $\mathcal{J}(\mathbf{X};\mathbf{u}):\mathbb{R}^{k}\rightarrow\mathbb{R}^{n}$ is the Jacobian of the function $\mathcal{H}(\mathbf{X};\mathbf{u})$. Recall that $\mathcal{S}^{k-1}$ denotes the unit sphere on $\mathbb{R}^{k}$, then we define
$\beta=\sup\{\|\mathcal{H}(\mathbf{X};\mathbf{w})\|_{2}:\mathbf{w}\in\mathcal{S}^{k-1}\},\text{ and }\alpha=\sup\{\|\mathcal{J}(\mathbf{X};\mathbf{w})\|_{2}:\mathbf{w}\in\mathcal{S}^{k-1}\}.$ (21)
We will start by establishing some auxiliary lemmata. The following lemma follows from $L$-homogeneity of $\mathcal{N}(\mathbf{u})$ and Lemma A.2.
Lemma B.1.
For any $\mathbf{u}\in\mathbb{R}^{k}$ and $c\geq 0$, $\nabla\mathcal{N}(\mathbf{u})^{\top}\mathbf{u}=L\mathcal{N}(\mathbf{u})$ and $\nabla\mathcal{N}(c\mathbf{u})=c^{L-1}\nabla\mathcal{N}(\mathbf{u})$, and $\mathcal{J}(\mathbf{X};\mathbf{u})\mathbf{u}=L\mathcal{H}(\mathbf{X};\mathbf{u})$ and $\mathcal{J}(\mathbf{X};c\mathbf{u})=c^{L-1}\mathcal{J}(\mathbf{X};\mathbf{u}).$
We next define $\tilde{\mathcal{N}}(\mathbf{u})$, which plays an integral part in the proof, and its gradient.
Lemma B.2.
For any nonzero $\mathbf{u}\in\mathbb{R}^{k}$ we define $\tilde{\mathcal{N}}(\mathbf{u})=\mathcal{N}(\mathbf{u})/\|\mathbf{u}\|_{2}^{L}$, then,
$\nabla\tilde{\mathcal{N}}(\mathbf{u})=\frac{\nabla\mathcal{N}(\mathbf{u})}{\|\mathbf{u}\|_{2}^{L}}-\frac{L\mathcal{N}(\mathbf{u})\mathbf{u}}{\|\mathbf{u}\|_{2}^{L+2}}=\left(\mathbf{I}-\frac{\mathbf{u}\mathbf{u}^{\top}}{\|\mathbf{u}\|_{2}^{2}}\right)\frac{\nabla\mathcal{N}(\mathbf{u})}{\|\mathbf{u}\|_{2}^{L}}.$
Proof.
The first equality follows by differentiating $\tilde{\mathcal{N}}(\mathbf{u})$ with respect to $\mathbf{u}$; using Lemma B.1 we get the second equality. ∎
We next describe the conditions for first-order KKT point of the NCF.
Lemma B.3.
If a vector $\mathbf{u}^{*}\in\mathbb{R}^{k\times 1}$ is a first-order KKT point of
$\max_{\|\mathbf{u}\|_{2}^{2}=1}\mathcal{N}(\mathbf{u})=\mathbf{z}^{\top}\mathcal{H}(\mathbf{X};\mathbf{u}),$ (22)
then
$\nabla\mathcal{N}(\mathbf{u}^{*})=\mathcal{J}(\mathbf{X};\mathbf{u}^{*})^{\top}\mathbf{z}=\lambda^{*}\mathbf{u}^{*},\|\mathbf{u}^{*}\|_{2}^{2}=1,$ (23)
where $\lambda^{*}\in\mathbb{R}$ is the Lagrange multiplier. Also, $L\mathcal{N}(\mathbf{u}^{*})=\lambda^{*}$, which implies that $\lambda^{*}\geq 0$ for a non-negative KKT point.
Proof.
For (22), the Lagrangian is
$L(\mathbf{u},\lambda)=\mathcal{N}(\mathbf{u})+\lambda(\|\mathbf{u}\|_{2}^{2}-1).$
Hence, if $\mathbf{u}^{*}$ is a first-order KKT point then $\|\mathbf{u}^{*}\|_{2}^{2}=1$, and for some $\overline{\lambda}$ we have
$\nabla\mathcal{N}(\mathbf{u}^{*})+2\overline{\lambda}\mathbf{u}^{*}=\mathbf{0},$
which implies
$\mathcal{J}(\mathbf{X};\mathbf{u}^{*})^{\top}\mathbf{z}+2\overline{\lambda}\mathbf{u}^{*}=\mathbf{0}.$
We get (23) by choosing $\lambda^{*}=-2\overline{\lambda}$. Further, by Lemma B.1, we have
$\lambda^{*}=\lambda^{*}\|\mathbf{u}^{*}\|_{2}^{2}={\mathbf{u}^{*}}^{\top}\nabla\mathcal{N}(\mathbf{u}^{*})=L\mathcal{N}(\mathbf{u}^{*}).$
Lemma B.4.
For any $t\geq 0$, we have
$\frac{d\mathcal{N}(\mathbf{u})}{dt}=\|\nabla\mathcal{N}(\mathbf{u}(t))\|_{2}^{2}=\|\dot{\mathbf{u}}(t)\|_{2}^{2}.$ (24)
Proof.
Using $\dot{\mathbf{u}}=\nabla\mathcal{N}(\mathbf{u}(t))$, we have
$\frac{d\mathcal{N}(\mathbf{u})}{dt}=\nabla\mathcal{N}(\mathbf{u}(t))^{\top}\dot{\mathbf{u}}=\|\nabla\mathcal{N}(\mathbf{u}(t))\|_{2}^{2}=\|\dot{\mathbf{u}}(t)\|_{2}^{2}.$
Lemma B.5.
For any $t_{0}\geq 0$, if $\mathcal{N}(\mathbf{u}(t_{0}))>0,$ then, for all $t\geq t_{0}$, we have $\mathcal{N}(\mathbf{u}(t))\geq\mathcal{N}(\mathbf{u}(t_{0}))>0$ and $\|\mathbf{u}(t)\|_{2}\geq\|\mathbf{u}(t_{0})\|_{2}$.
Proof.
Using Lemma B.4, we have
[…]
Therefore, for $t\geq t_{0}$, $\mathcal{N}(\mathbf{u}(t))\geq\mathcal{N}(\mathbf{u}(t_{0}))>0$. The second claim is true since for all $t\geq 0$
$\frac{d{\|\mathbf{u}\|_{2}^{2}}}{dt}=2\mathbf{u}^{\top}\dot{\mathbf{u}}=2\mathbf{u}^{\top}\nabla\mathcal{N}(\mathbf{u})=2L\mathcal{N}(\mathbf{u}),$
which implies
[…]
We now turn to proving Lemma 5.2.
Proof of Lemma 5.2.
We begin by showing that either the limit of $\mathbf{u}(t)/\|\mathbf{u}(t)\|_{2}$ exists or $\mathbf{u}(t)$ converges to $\mathbf{0}$. For this, we consider two cases. Case 1: $\mathcal{N}(\mathbf{u}(0))>0$.
In this case, we first show that $\mathbf{u}(t)$ becomes unbounded at some finite time. Suppose for the sake of contradiction $\|\mathbf{u}(t)\|_{2}<\infty$ for all finite $t\geq 0$. Let $\mathcal{N}(\mathbf{u}(0))=\gamma>0,$ thus $\|\mathbf{u}(0)\|_{2}>0$. Using Lemma B.5, for all $t\geq 0$,
$\|\mathbf{u}(t)\|_{2}\geq\|\mathbf{u}(0)\|_{2}>0.$
Hence, $\tilde{\mathcal{N}}(\mathbf{u}(t))$ (which we recall is equal to $\mathcal{N}(\mathbf{u}(t))/\|\mathbf{u}(t)\|_{2}^{L}$) is defined for all $t\geq 0$. Now, by Lemma B.4, for all $t\geq 0$, we have
$\frac{d{{\mathcal{N}}(\mathbf{u})}}{dt}=\|\dot{\mathbf{u}}\|_{2}^{2},\text{ and }\dot{\mathbf{u}}={\nabla}\mathcal{N}(\mathbf{u}).$
Therefore, using Lemma B.2, for all $t\geq 0$,
$\frac{d{\tilde{\mathcal{N}}(\mathbf{u})}}{dt}=\dot{\mathbf{u}}^{\top}\nabla\tilde{\mathcal{N}}(\mathbf{u})=\frac{{\nabla}\mathcal{N}(\mathbf{u})^{\top}}{\|\mathbf{u}\|_{2}^{L}}\left(\mathbf{I}-\frac{\mathbf{u}\mathbf{u}^{T}}{\|\mathbf{u}\|_{2}^{2}}\right){\nabla}\mathcal{N}(\mathbf{u})\geq 0.$ (25)
For all $t_{2}\geq t_{1}\geq 0$, integrating the above equality on both sides from $t_{1}$ to $t_{2}$ we get
[…]
Thus, $\tilde{\mathcal{N}}(\mathbf{u}(t))$ is an increasing function. Now if we define $\tilde{\mathcal{N}}(\mathbf{u}(0))=\tilde{\gamma}$, then for any $t\geq 0$ we have
$\tilde{{\mathcal{N}}}(\mathbf{u}(t))\geq\tilde{\gamma}>0,$
which implies
$\mathcal{N}(\mathbf{u}(t))\geq\tilde{\gamma}\|\mathbf{u}(t)\|_{2}^{L},\forall t\geq 0.$
From the above inequality, we have
$\frac{1}{2}\frac{d{\|\mathbf{u}\|_{2}^{2}}}{dt}=\mathbf{u}^{\top}\dot{\mathbf{u}}=L\mathcal{N}(\mathbf{u}(t))\geq L\tilde{\gamma}\|\mathbf{u}(t)\|_{2}^{L},\forall t\geq 0.$ (27)
Taking $\|\mathbf{u}(t)\|_{2}^{L}$ to the LHS and integrating both sides from 00 to $t$, we get
$\frac{1}{{L}/{2}-1}\left(\frac{1}{\|\mathbf{u}(0)\|_{2}^{L/2-1}}-\frac{1}{\|\mathbf{u}(t)\|_{2}^{L/2-1}}\right)\geq 2L\tilde{\gamma}t,\forall t\geq 0.$
Simplifying the above equation gives us
$\|\mathbf{u}(t)\|_{2}^{L/2-1}\geq\frac{\|\mathbf{u}(0)\|_{2}^{L/2-1}}{1-t\tilde{\gamma}L(L-2)\|\mathbf{u}(0)\|_{2}^{L/2-1}},\forall t\geq 0.$
The above inequality indicates that $\|\mathbf{u}(t)\|_{2}$ is not bounded for all finite $t\geq 0$, which leads to a contradiction. Henceforth, we choose $T^{*}$ to denote the time when $\mathbf{u}(t)$ becomes infinity. Formally, we assume that
$\lim_{t\rightarrow T^{*}}\|\mathbf{u}(t)\|_{2}=\infty,\text{ and }\|\mathbf{u}(t)\|_{2}<\infty,\text{ for all }t\in[0,T^{*}).$
We now proceed to prove that $\lim_{t\rightarrow T^{*}}\mathbf{u}(t)/\|\mathbf{u}(t)\|_{2}$ exists by establishing that the length of the curve swept by $\mathbf{u}(t)/\|\mathbf{u}(t)\|_{2}$, which is defined as
[…]
is of finite length. This proof technique is inspired from [31], and a similar technique was also employed in [1]. From (25) and (26), we know $\tilde{\mathcal{N}}(\mathbf{u}(t))$ is an increasing function in $[0,T^{*})$. Further, since $\mathcal{N}(\mathbf{u})=\mathbf{z}^{\top}\mathcal{H}(\mathbf{X};\mathbf{u})\leq\beta\|\mathbf{z}\|_{2}\|\mathbf{u}\|_{2}^{L}$, where the inequality follows from (21), we have that $\tilde{\mathcal{N}}(\mathbf{u})$ is bounded from above. Thus, using monotone convergence theorem, we have $\lim_{t\rightarrow T^{*}}\tilde{\mathcal{N}}(\mathbf{u}(t))$ exists.
Now, for all $t\in[0,T^{*})$, we know
$\frac{d}{dt}\left(\frac{\mathbf{u}}{\|\mathbf{u}\|_{2}}\right)=\left(\mathbf{I}-\frac{\mathbf{u}\mathbf{u}^{T}}{\|\mathbf{u}\|_{2}^{2}}\right)\frac{\dot{\mathbf{u}}}{\|\mathbf{u}\|_{2}}=\left(\mathbf{I}-\frac{\mathbf{u}\mathbf{u}^{T}}{\|\mathbf{u}\|_{2}^{2}}\right)\frac{{\nabla}\mathcal{N}(\mathbf{u})}{\|\mathbf{u}\|_{2}}.$ (28)
Therefore,
$\left\|\frac{d}{dt}\left(\frac{\mathbf{u}}{\|\mathbf{u}\|_{2}}\right)\right\|_{2}=\left\|\left(\mathbf{I}-\frac{\mathbf{u}\mathbf{u}^{T}}{\|\mathbf{u}\|_{2}^{2}}\right){\nabla}\mathcal{N}(\mathbf{u})\right\|_{2}\frac{1}{\|\mathbf{u}\|_{2}}.$ (29)
Now, suppose $\lim_{t\rightarrow T^{*}}\tilde{\mathcal{N}}(\mathbf{u}(t))=\overline{f}$. If $\tilde{\mathcal{N}}(\mathbf{u}(t))$ converges to $\overline{f}$ before $T^{*}$, i.e., $\tilde{\mathcal{N}}(\mathbf{u}(T))=\overline{f}$ for some $T<T^{*}$. Then, for all $t\in[T,T^{*})$,
$\frac{d{\tilde{\mathcal{N}}(\mathbf{u})}}{dt}=0,$
which implies
$\left\|\left(\mathbf{I}-\frac{\mathbf{u}(t)\mathbf{u}(t)^{T}}{\|\mathbf{u}(t)\|_{2}^{2}}\right){\nabla}\mathcal{N}(\mathbf{u}(t))\right\|_{2}=0,\forall t\in[T,T^{*}).$
Hence, from (29), we have
$\frac{d}{dt}\left(\frac{\mathbf{u}}{\|\mathbf{u}\|_{2}}\right)=\mathbf{0},\forall t\in[T,T^{*}),$
which implies $\lim_{t\rightarrow T^{*}}\frac{\mathbf{u}(t)}{\|\mathbf{u}(t)\|_{2}}$ exists and is equal to $\frac{\mathbf{u}(T)}{\|\mathbf{u}(T)\|_{2}}$. Thus, we may assume $\overline{f}-\tilde{\mathcal{N}}(\mathbf{u}(t))>0$, for all $t\in[0,T^{*})$. We define $g(\mathbf{u})=\overline{f}-\tilde{\mathcal{N}}(\mathbf{u})$. Then, since
$\|{\nabla}_{r}\tilde{\mathcal{N}}(\mathbf{u})\|_{2}=0,$
we have that
$\|{\nabla}_{\perp}g(\mathbf{u})\|_{2}\geq 0=\|\mathbf{u}\|_{2}\|\overline{\nabla}_{r}g(\mathbf{u})\|_{2}.$
Hence, from A.1, there exists a $\nu>0$ and a desingularizing function $\Psi(.)$ defined on $[0,\nu)$ such that if $\|\mathbf{u}\|_{2}>1$ and $g(\mathbf{u})<\nu$, then
$1\geq\Psi^{\prime}(g(\mathbf{u}))\|\mathbf{u}\|_{2}\|{\nabla}g(\mathbf{u})\|_{2}=\Psi^{\prime}(\overline{f}-\tilde{\mathcal{N}}(\mathbf{u}))\|\mathbf{u}\|_{2}\|{\nabla}\tilde{\mathcal{N}}(\mathbf{u})\|_{2}.$ (30)
Since $\lim_{t\rightarrow T^{*}}\tilde{\mathcal{N}}(\mathbf{u}(t))=\overline{f}$, and $\lim_{t\rightarrow T^{*}}\|\mathbf{u}(t)\|_{2}=\infty$, we may choose $T$ large enough such that $\|\mathbf{u}(t)\|_{2}>1$, and $g(\mathbf{u}(t))<\nu$, for all $t\in[T,T^{*})$. Hence, for all $t\in[T,T^{*})$, we have
$\frac{d{\tilde{\mathcal{N}}(\mathbf{u})}}{dt}$ $=\frac{{\nabla}\mathcal{N}(\mathbf{u})^{\top}}{\|\mathbf{u}\|_{2}^{L}}\left(\mathbf{I}-\frac{\mathbf{u}\mathbf{u}^{T}}{\|\mathbf{u}\|_{2}^{2}}\right){\nabla}\mathcal{N}(\mathbf{u})$
$=\left\|\left(\mathbf{I}-\frac{\mathbf{u}\mathbf{u}^{T}}{\|\mathbf{u}\|_{2}^{2}}\right)\frac{{\nabla}\mathcal{N}(\mathbf{u})}{\|\mathbf{u}\|_{2}^{L-1}}\right\|_{2}\left\|\frac{d}{dt}\left(\frac{\mathbf{u}}{\|\mathbf{u}\|_{2}}\right)\right\|_{2}$
$=\|\mathbf{u}\|_{2}\left\|{\nabla}\tilde{\mathcal{N}}(\mathbf{u})\right\|_{2}\left\|\frac{d}{dt}\left(\frac{\mathbf{u}}{\|\mathbf{u}\|_{2}}\right)\right\|_{2}\leq\frac{1}{\Psi^{\prime}(\overline{f}-\tilde{\mathcal{N}}(\mathbf{u}))}\left\|\frac{d}{dt}\left(\frac{\mathbf{u}}{\|\mathbf{u}\|_{2}}\right)\right\|_{2}.$
In the above chain of equalities and inequalities, we used Lemma B.2 in the third equality, and the inequality follows from (30). Now, since $\overline{f}-\tilde{\mathcal{N}}(\mathbf{u}(t))\in[0,\nu),$ for all $t\in[T,T^{*})$, and $\Psi^{\prime}>0$ on $[0,\nu)$, we can take $\Psi^{\prime}(\overline{f}-\tilde{\mathcal{N}}(\mathbf{u}(t)))$ to the left hand side to get
$\left\|\frac{d}{dt}\left(\frac{\mathbf{u}}{\|\mathbf{u}\|_{2}}\right)\right\|_{2}$ $\leq-\frac{d\Psi(\overline{f}-\tilde{\mathcal{N}}(\mathbf{u}))}{dt},\forall t\in[T,T^{*}).$
Next, integrating the above inequality on both the sides from $T$ to any $t_{1}\in[T,T^{*})$, we have
$\int_{T}^{t_{1}}\left\|\frac{d}{dt}\left(\frac{\mathbf{u}}{\|\mathbf{u}\|_{2}}\right)\right\|_{2}dt\leq\Psi(\overline{f}-\tilde{\mathcal{N}}(\mathbf{u}(T))-\Psi(\overline{f}-\tilde{f}(\mathbf{u}(t_{1}))\leq\Psi(\overline{f}-\tilde{\mathcal{N}}(\mathbf{u}(T)))<\infty.$
From the above inequality, we further have
[…]
$\leq\int_{0}^{T}\left\|\frac{d}{dt}\left(\frac{\mathbf{u}}{\|\mathbf{u}\|_{2}}\right)\right\|_{2}dt+\Psi(\overline{f}-\tilde{\mathcal{N}}(\mathbf{u}(T))<\infty,$
which completes the proof. Case 2: $\mathcal{N}(\mathbf{u}(0))\leq 0$.
In this case, we may further assume that $\mathcal{N}(\mathbf{u}(t))\leq 0,$ for all $t\geq 0$, since if for some $\overline{t}$, $\mathcal{N}(\mathbf{u}(\overline{t}))>0,$ then we can choose $\overline{t}$ be the new starting time and use the proof for Case 1 to show that $\mathbf{u}(t)$ becomes unbounded at some finite time and also converges in direction. Therefore, we assume $\mathcal{N}(\mathbf{u}(t))\leq 0,$ for all $t\geq 0$. Now, since
$\frac{1}{2}\frac{d{\|\mathbf{u}\|_{2}^{2}}}{dt}=\mathbf{u}^{\top}\dot{\mathbf{u}}=L\mathcal{N}(\mathbf{u}(t))\leq 0,\forall t\geq 0,$
we have that $\|\mathbf{u}(t)\|_{2}$ decreases with time, and $\lim_{t\rightarrow\infty}\|\mathbf{u}(t)\|_{2}$ exists. Now, if $\lim_{t\rightarrow\infty}\|\mathbf{u}(t)\|_{2}=0$, then we have $\lim_{t\rightarrow\infty}\mathbf{u}(t)=\mathbf{0}$ and our proof is complete. Else, suppose $\lim_{t\rightarrow\infty}\|\mathbf{u}(t)\|_{2}=\eta>0$. Then, we have $\|\mathbf{u}(t)\|_{2}\geq\eta$, for all $t\geq 0$, since $\|\mathbf{u}(t)\|_{2}$ decreases with time. We next show that $\lim_{t\rightarrow\infty}\mathcal{N}(\mathbf{u}(t))=0$. From Lemma B.4, we know
$\frac{d{\mathcal{N}(\mathbf{u})}}{dt}=\|\dot{\mathbf{u}}\|_{2}^{2},\forall t\geq 0.$
Thus, $\mathcal{N}(\mathbf{u}(t))$ increases with time. Since we have assumed $\mathcal{N}(\mathbf{u}(t))\leq 0$ for all $t\geq 0$, by monotone convergence we have $\lim_{t\rightarrow\infty}\mathcal{N}(\mathbf{u}(t))$ exists. Now, to show $\lim_{t\rightarrow\infty}\mathcal{N}(\mathbf{u}(t))=0$, we assume for the sake of contradiction that $\lim_{t\rightarrow\infty}\mathcal{N}(\mathbf{u}(t))=-\gamma<0$. Then, for all $t\geq 0$, we have $\mathcal{N}(\mathbf{u}(t))\geq-\gamma$, since $\mathcal{N}(\mathbf{u}(t))$ is an increasing function of time. Thus,
$\frac{1}{2}\frac{d{\|\mathbf{u}\|_{2}^{2}}}{dt}=\mathbf{u}^{\top}\dot{\mathbf{u}}=L\mathcal{N}(\mathbf{u}(t))\leq-L\gamma,\forall t\geq 0.$
From the above equation we get that $\|\mathbf{u}(t)\|_{2}$ would become smaller than $\eta$ after some finite amount of time has elapsed. Since $\|\mathbf{u}(t)\|_{2}\geq\eta$, for all $t\geq 0$, this leads to a contradiction. Therefore, $\lim_{t\rightarrow\infty}\mathcal{N}(\mathbf{u}(t))=0$. This also implies $\lim_{t\rightarrow\infty}\tilde{\mathcal{N}}(\mathbf{u}(t))=0$. We next prove that $\lim_{t\rightarrow\infty}\mathbf{u}(t)/\|\mathbf{u}(t)\|_{2}$ exists. We first define $\hat{\mathbf{u}}(t)=2\mathbf{u}(t)/\eta$. Now, if $\hat{\mathbf{u}}(t)$ converges in direction, then ${\mathbf{u}}(t)$ also converges in direction. To prove $\hat{\mathbf{u}}(t)$ converges in direction we can follow the same approach as in Case 1, specifically from (29) onward. This transformation is essential because, recall, to use Lemma A.1 in Case 1 we had to choose $T$ large enough such that $\|\mathbf{u}(t)\|_{2}>1$, for all $t\geq T$. Here, $\|\mathbf{u}(t)\|_{2}$ may never be greater than $1$, but $\|\hat{\mathbf{u}}(t)\|_{2}\geq 2$, for all $t\geq 0$. So far we have established that either $\frac{\mathbf{u}(t)}{\|\mathbf{u}(t)\|_{2}}$ converges, or $\mathbf{u}(t)$ converges to $\mathbf{0}$. We next show that if $\frac{\mathbf{u}(t)}{\|\mathbf{u}(t)\|_{2}}$ converges to $\mathbf{u}^{*}$, then $\mathbf{u}^{*}$ must be a non-negative KKT point of the constrained NCF. From the proof so far we already know that $\mathcal{N}(\mathbf{u}^{*})=\tilde{\mathcal{N}}(\mathbf{u}^{*})\geq 0.$ Therefore, from Lemma B.3, we only need to show
$L\mathcal{N}(\mathbf{u}^{*})\mathbf{u}^{*}=\nabla\mathcal{N}(\mathbf{u}^{*}),$ (31)
For the sake of contradiction assume that there exists some $\gamma>0$ such that
$\left\|\nabla\mathcal{N}(\mathbf{u}^{*})-L\mathcal{N}(\mathbf{u}^{*})\mathbf{u}^{*}\right\|_{2}\geq\gamma.$ (32)
For any $\epsilon>0$, we define the set $\mathbf{u}_{\epsilon}=\{\mathbf{u}:\|\mathbf{u}-\mathbf{u}^{*}\|_{2}\leq\epsilon\}$. Since $\mathcal{N}(\mathbf{u})$ has locally Lipschitz gradient, given $\gamma$, we can choose sufficiently small $\epsilon\in(0,1)$ such that for all $\mathbf{u}\in\mathbf{u}_{\epsilon}$, we have
$\|{\nabla}\mathcal{N}(\mathbf{u})-\nabla\mathcal{N}(\mathbf{u}^{*})\|_{2}\leq\gamma/4.$ (33)
Since $\frac{\mathbf{u}(t)}{\|\mathbf{u}(t)\|_{2}}$ converges to $\mathbf{u}^{*}$ and $\mathcal{N}(\mathbf{u})$ is continuous, we choose $T$ large enough such that for all $t\geq T$,
$\left\|\frac{\mathbf{u}(t)}{\|\mathbf{u}(t)\|_{2}}-\mathbf{u}^{*}\right\|_{2}\leq\epsilon,\left\|\frac{L\mathbf{u}(t)}{\|\mathbf{u}(t)\|_{2}}\mathcal{N}\left(\frac{\mathbf{u}(t)}{\|\mathbf{u}(t)\|_{2}}\right)-L\mathbf{u}^{*}\mathcal{N}(\mathbf{u}^{*})\right\|\leq\gamma/4.$ (34)
From the two cases discussed above we know that if $\frac{\mathbf{u}(t)}{\|\mathbf{u}(t)\|_{2}}$ converges, then $\|\mathbf{u}(t)\|_{2}$ either goes to infinity or remains away from 00. Thus, we may also assume that $T$ is large enough such that for some $\eta>0$, we have
$\|\mathbf{u}(t)\|_{2}\geq\eta,\forall t\geq T.$
Now, for all non-zero $\mathbf{u}\in\mathbb{R}^{k}$ we have
$\left\|\left(\mathbf{I}-\frac{\mathbf{u}\mathbf{u}^{\top}}{\|\mathbf{u}\|_{2}^{2}}\right)\frac{{\nabla}\mathcal{N}(\mathbf{u})}{\|\mathbf{u}\|_{2}^{L-1}}\right\|_{2}=\left\|\frac{{\nabla}\mathcal{N}(\mathbf{u})}{\|\mathbf{u}\|_{2}^{L-1}}-\frac{L\mathbf{u}\mathcal{N}(\mathbf{u})}{\|\mathbf{u}\|_{2}^{L+1}}\right\|_{2}$
$=$ $\left\|{\nabla}\mathcal{N}\left(\frac{\mathbf{u}}{\|\mathbf{u}\|_{2}}\right)-\frac{L\mathbf{u}}{\|\mathbf{u}\|_{2}}\mathcal{N}\left(\frac{\mathbf{u}}{\|\mathbf{u}\|_{2}}\right)\right\|_{2}$
$\geq$ $\left\|\nabla\mathcal{N}(\mathbf{u}^{*})-\frac{L\mathbf{u}}{\|\mathbf{u}\|_{2}}\mathcal{N}\left(\frac{\mathbf{u}}{\|\mathbf{u}\|_{2}}\right)\right\|_{2}-\left\|{\nabla}\mathcal{N}\left(\frac{\mathbf{u}}{\|\mathbf{u}\|_{2}}\right)-\nabla\mathcal{N}(\mathbf{u}^{*})\right\|_{2}$
$\geq$ $\left\|\nabla\mathcal{N}(\mathbf{u}^{*})-{L\mathbf{u}^{*}}\mathcal{N}\left({\mathbf{u}^{*}}\right)\right\|_{2}-\left\|{\nabla}\mathcal{N}\left(\frac{\mathbf{u}}{\|\mathbf{u}\|_{2}}\right)-\nabla\mathcal{N}(\mathbf{u}^{*})\right\|_{2}-\left\|{L\mathbf{u}^{*}}\mathcal{N}\left({\mathbf{u}^{*}}\right)-\frac{L\mathbf{u}}{\|\mathbf{u}\|_{2}}\mathcal{N}\left(\frac{\mathbf{u}}{\|\mathbf{u}\|_{2}}\right)\right\|_{2},$
where in the second equality we used $(L-1)$-homogeneity of $\nabla\mathcal{N}(\mathbf{u})$ and $L$-homogeneity of $\mathcal{N}(\mathbf{u})$. The inequalities make repeated use of triangle inequality of norms. Hence, using (32), (33) and (34), for all $t\geq T$, we have
$\left\|\left(\mathbf{I}-\frac{\mathbf{u}(t)\mathbf{u}(t)^{\top}}{\|\mathbf{u}(t)\|_{2}^{2}}\right)\frac{{\nabla}\mathcal{N}(\mathbf{u}(t))}{\|\mathbf{u}(t)\|_{2}^{L-1}}\right\|_{2}\geq\gamma/2.$
Using the above inequality and (25) we have
$\frac{d}{dt}\left(\mathcal{N}\left(\frac{\mathbf{u}(t)}{\|\mathbf{u}(t)\|_{2}}\right)\right)=\left\|\left(\mathbf{I}-\frac{\mathbf{u}(t)\mathbf{u}(t)^{\top}}{\|\mathbf{u}(t)\|_{2}^{2}}\right)\frac{{\nabla}\mathcal{N}(\mathbf{u}(t))}{\|\mathbf{u}(t)\|_{2}^{L/2}}\right\|_{2}^{2}\geq\gamma^{2}\|\mathbf{u}(t)\|_{2}^{L-2}/4\geq\gamma^{2}\eta^{L-2}/4.$
However, since the limit of $\mathcal{N}\left(\frac{\mathbf{u}(t)}{\|\mathbf{u}(t)\|_{2}}\right)$ exists, the above inequality leads to a contradiction, which completes our proof. ∎
Before proceeding to proof of Theorem 5.1, we state a useful lemma.
Lemma B.6.
Let $\mathbf{u}_{0}$ be a fixed vector and let $\mathbf{u}(t)$ be the solution of
$\frac{d{\mathbf{u}}}{dt}=\nabla\mathcal{N}(\mathbf{u})=\mathcal{J}(\mathbf{X};\mathbf{u})^{\top}\mathbf{z},{\mathbf{u}}(0)=\mathbf{u}_{0}.$ (35)
If $\lim_{t\rightarrow\infty}\mathbf{u}(t)\neq\mathbf{0},$ then there exists $\eta>0$ and $T\geq 0$ such that $\|\mathbf{u}(t)\|_{2}\geq\eta$, for all $t\geq T$. Further, for any $\epsilon\in(0,1)$ there exists $T_{\epsilon}\geq T$ and $B_{\epsilon}$ such that
$\frac{\mathbf{u}(T_{\epsilon})^{\top}{\mathbf{u}^{*}}}{\|\mathbf{u}(T_{\epsilon})\|_{2}}\geq 1-\epsilon,\text{ and }\|\mathbf{u}(t)\|_{2}\leq B_{\epsilon},\forall t\leq T_{\epsilon},$ (36)
where $\mathbf{u}^{*}$ is a non-negative KKT point of
$\max_{\|\mathbf{u}\|_{2}^{2}=1}\mathcal{N}(\mathbf{u})=\mathbf{z}^{\top}\mathcal{H}(\mathbf{X};\mathbf{u}).$ (37)
Proof.
Similar to the proof of Lemma 5.2, we consider two cases. Case 1: $\mathcal{N}(\mathbf{u}(0))>0$.
Let $\mathcal{N}(\mathbf{u}(0))=\gamma>0,$ thus $\|\mathbf{u}(0)\|_{2}>0$. From Lemma B.5, $\|\mathbf{u}(t)\|_{2}\geq\|\mathbf{u}(0)\|_{2}$, for all $t\geq 0$, which implies we can choose $\eta=\|\mathbf{u}(0)\|_{2}$ and $T=0$. Next, from Lemma 5.2 and its proof we know that for some finite $T^{*}$, $\lim_{t\rightarrow T^{*}}\frac{\mathbf{u}(t)}{\|\mathbf{u}(t)\|_{2}}=\mathbf{u}^{*}$, where $\mathbf{u}^{*}$ is a non-negative KKT point of the constrained NCF (37). Thus, for any $\epsilon\in(0,1)$, we can choose $T_{\epsilon}\geq T=0$, and large enough but less than $T^{*}$ such that
$\frac{\mathbf{u}(T_{\epsilon})^{\top}{\mathbf{u}^{*}}}{\|\mathbf{u}(T_{\epsilon})\|_{2}}\geq 1-\epsilon.$ (38)
Since for all $t\in(0,T^{*})$, $\mathbf{u}(t)$ is finite, there exists $B_{\epsilon}$ such that $\|\mathbf{u}(T_{\epsilon})\|_{2}\leq B_{\epsilon}.$ We note that as $\epsilon$ gets smaller, we may have to choose $T_{\epsilon}$ closer to $T^{*}$. This implies that, since $\lim_{t\rightarrow T^{*}}\|\mathbf{u}(t)\|_{2}=\infty,$ $B_{\epsilon}$ will increase. However, for a fixed $\epsilon$, both $T_{\epsilon}$ and $B_{\epsilon}$ will be fixed.
Case 2: $\mathcal{N}(\mathbf{u}(0))\leq 0$.
In this case, we can also assume that $\mathcal{N}(\mathbf{u}(t))\leq 0,$ for all $t\geq 0$, since if $\mathcal{N}(\mathbf{u}(\overline{t}))>0$, for some $\overline{t}$, then from Lemma B.5, we know $\|\mathbf{u}(t)\|_{2}\geq\|\mathbf{u}(\overline{t})\|_{2}$, for all $t\geq\overline{t}$. Moreover, since $\mathcal{N}(\mathbf{u}(\overline{t}))>0$ implies $\|\mathbf{u}(\overline{t})\|_{2}>0$, we can choose $\eta=\|\mathbf{u}(\overline{t})\|_{2}$ and $T=\overline{t}$. Also, in this case, from the proof of Lemma 5.2, we know that for some finite $T^{*}$, $\lim_{t\rightarrow T^{*}}\frac{\mathbf{u}(t)}{\|\mathbf{u}(t)\|_{2}}=\mathbf{u}^{*}$, where $\mathbf{u}^{*}$ is a non-negative KKT point of the constrained NCF (37). Thus, similar to Case 1, for any $\epsilon\in(0,1)$ we can choose $T_{\epsilon}\geq\overline{t}$ and $B_{\epsilon}$ as desired. Hence, suppose $\mathcal{N}(\mathbf{u}(t))\leq 0,$ for all $t\geq 0$. Then, since
$\frac{1}{2}\frac{d{\|\mathbf{u}\|_{2}^{2}}}{dt}=\mathbf{u}^{\top}\dot{\mathbf{u}}=L\mathcal{N}(\mathbf{u}(t))\leq 0,\forall t\geq 0,$
we have that $\|\mathbf{u}(t)\|_{2}$ is a decreasing function with time, and hence, $\lim_{t\rightarrow\infty}\|\mathbf{u}(t)\|_{2}$ exists. Also,
$\|\mathbf{u}(t)\|_{2}\geq\lim_{t\rightarrow\infty}\|\mathbf{u}(t)\|_{2},\text{for all }t\geq 0.$
Since we have assumed $\lim_{t\rightarrow\infty}\mathbf{u}(t)\neq\mathbf{0},$ we know $\lim_{t\rightarrow\infty}\|\mathbf{u}(t)\|_{2}>0$, and we can choose $\eta=\lim_{t\rightarrow\infty}\|\mathbf{u}(t)\|_{2}$ and $T=0$. Further, since $\lim_{t\rightarrow\infty}\frac{\mathbf{u}(t)}{\|\mathbf{u}(t)\|_{2}}=\mathbf{u}^{*}$, where $\mathbf{u}^{*}$ is a non-negative KKT point of the constrained NCF (37), for any $\epsilon\in(0,1)$, we can choose $T_{\epsilon}$ large enough such that
$\frac{\mathbf{u}(T_{\epsilon})^{\top}{\mathbf{u}^{*}}}{\|\mathbf{u}(T_{\epsilon})\|_{2}}\geq 1-\epsilon.$ (39)
Also, since $\|\mathbf{u}(t)\|_{2}$ is a decreasing function with time, we have $\|\mathbf{u}(t)\|_{2}\in[\eta,\|\mathbf{u}(0)\|_{2}]$, for all $t\geq 0$. Thus, we can choose $B_{\epsilon}=\|\mathbf{u}(0)\|_{2}$. ∎
B.2 Proof of Theorem 5.1
Proof.
Let $\mathbf{w}_{0}$ be a fixed unit norm vector. Suppose $\mathbf{u}(t)$ is the solution of
$\frac{d\mathbf{u}}{dt}=\nabla\mathcal{N}_{\mathbf{y},\mathcal{H}}(\mathbf{u})=\mathcal{J}\left(\mathbf{X};\mathbf{u}\right)^{\top}\mathbf{y},\mathbf{u}(0)=\mathbf{w}_{0}.$ (40)
From Lemma 5.2, we know that there are two possibilities. Either $\mathbf{u}(t)$ converges to $\mathbf{0}$, or $\mathbf{u}(t)$ converges in direction to a non-negative KKT point of the constrained NCF. If $\mathbf{u}(t)$ converges to $\mathbf{0}$, then we define $\eta=2.$ Then, for any $\epsilon\in(0,\eta/2)$, we define $B_{\epsilon}=\epsilon$ and choose $T_{\epsilon}$ large enough such that
$\|\mathbf{u}(T_{\epsilon})\|_{2}\leq B_{\epsilon}=\epsilon.$ (41)
Else, if $\mathbf{u}(t)$ does not converge to $\mathbf{0}$, then from Lemma B.6, there exists $\tilde{\eta}>0$ and $T$ such that $\|\mathbf{u}(t)\|_{2}\geq\tilde{\eta}$, for all $t\geq T$. We further define $\eta=\min(\tilde{\eta},1)$. Then, from Lemma B.6, for any $\epsilon\in(0,\eta/2)$ there exists $T_{\epsilon}\geq T$ and $B_{\epsilon}$ such that
$\frac{\mathbf{u}(T_{\epsilon})^{\top}{\mathbf{u}^{*}}}{\|\mathbf{u}(T_{\epsilon})\|_{2}}\geq 1-\epsilon,\text{ and }\|\mathbf{u}(t)\|_{2}\leq B_{\epsilon},\forall t\leq T_{\epsilon},$ (42)
where $\mathbf{u}^{*}$ is a non-negative KKT point of
$\max_{\|\mathbf{u}\|_{2}^{2}=1}\mathcal{N}_{\mathbf{y},\mathcal{H}}(\mathbf{u})=\mathcal{H}\left(\mathbf{X};\mathbf{u}\right)^{\top}\mathbf{y}.$
Further, since $T_{\epsilon}\geq T$, we have
$\|\mathbf{u}(T_{\epsilon})\|_{2}\geq\tilde{\eta}\geq\eta.$ (43)
We next define other $\epsilon$ dependent parameters which are useful in our proof. Let $\tilde{B}_{\epsilon}=B_{\epsilon}+\epsilon$. Since $\mathcal{H}(\mathbf{x};\mathbf{w})$ has locally Lipschitz gradient, there exists $K_{\epsilon}>0$ such that if $\|\mathbf{w}_{1}\|_{2},\|\mathbf{w}_{2}\|_{2}\leq\tilde{B}_{\epsilon}$, then
$\|\mathcal{J}\left(\mathbf{X};\mathbf{w}_{1}\right)-\mathcal{J}\left(\mathbf{X};\mathbf{w}_{2}\right)\|_{2}\leq K_{\epsilon}\|\mathbf{w}_{1}-\mathbf{w}_{2}\|_{2}.$ (44)
We further define $C_{\epsilon}=\max(\beta\tilde{B}_{\epsilon}^{L},\|\mathbf{y}\|^{2}e^{\|\mathbf{y}\|\beta}\beta\tilde{B}_{\epsilon}^{L})$, and
$\bar{\delta}^{L}=\min\left(\frac{1}{\tilde{B}^{L}_{\epsilon}},\frac{\epsilon}{4\alpha\tilde{B}_{\epsilon}^{L-1}C_{\epsilon}T_{\epsilon}e^{2T_{\epsilon}K_{\epsilon}\|\mathbf{y}\|_{2}}}\right),$ (45)
where recall
$\beta=\sup\{\|\mathcal{H}(\mathbf{X};\mathbf{w})\|_{2}:\mathbf{w}\in\mathcal{S}^{k-1}\},\text{ and }\alpha=\sup\{\|\mathcal{J}(\mathbf{X};\mathbf{w})\|_{2}:\mathbf{w}\in\mathcal{S}^{k-1}\}.$ (46)
Now, for any $\delta\in(0,\bar{\delta})$, let $\mathbf{w}(t)$ be the solution of
$\dot{\mathbf{w}}(t)=-\nabla\mathcal{L}(\mathbf{w}(t)),\mathbf{w}(0)=\delta\mathbf{w}_{0},$ (47)
for square or logistic loss. Then, we can rewrite the above differential equation as
$\dot{\mathbf{w}}=-\mathcal{J}(\mathbf{X};\mathbf{w})^{\top}\mathbf{e},{\mathbf{w}}(0)=\delta\mathbf{w}_{0},$ (48)
where $\mathbf{e}$ is an $n-$dimensional vector such that ${e}_{i}:=\nabla_{\hat{y}}\ell\left(\mathcal{H}(\mathbf{x}_{i};\mathbf{w}),y_{i}\right).$ For square loss, $\nabla_{\hat{y}}\ell(\hat{y},y)=\hat{y}-y$, and for logistic loss, $\nabla_{\hat{y}}\ell(\hat{y},y)=-(2ye^{-\hat{y}y})/(1+e^{-\hat{y}y})$. We define $\xi(t):=\mathbf{e}(t)+\mathbf{y}$, then, for square loss
$\xi(t)=\mathcal{H}(\mathbf{X};\mathbf{w}(t)),$
and for logistic loss
$\xi_{i}(t)=y_{i}\left(\frac{1-e^{-y_{i}\mathcal{H}(\mathbf{x}_{i};\mathbf{w}(t))}}{1+e^{-y_{i}\mathcal{H}(\mathbf{x}_{i};\mathbf{w}(t))}}\right).$
Thus, the dynamics of $\mathbf{w}(t)$ can be written as
$\dot{\mathbf{w}}\in-\mathcal{J}(\mathbf{X};\mathbf{w}(t))^{\top}\mathbf{e}(t)=\mathcal{J}(\mathbf{X};\mathbf{w}(t))^{\top}(\mathbf{y}-\xi(t)).$ (49)
Now, define $\mathbf{s}(t)=\frac{1}{\delta}\mathbf{w}\left(\frac{t}{\delta^{L-2}}\right)$, then $\mathbf{s}(0)=\mathbf{w}_{0}$, and
$\frac{d\mathbf{s}}{dt}=\frac{d}{dt}\left(\frac{1}{\delta}\mathbf{w}\left(\frac{t}{\delta^{L-2}}\right)\right)$ $=\frac{1}{\delta^{L-1}}\dot{\mathbf{w}}\left(\frac{t}{\delta^{L-2}}\right)$
$=\frac{1}{\delta^{L-1}}\mathcal{J}\left(\mathbf{X};\mathbf{w}\left(\frac{t}{\delta^{L-2}}\right)\right)^{\top}\left(\mathbf{y}-\xi\left(\frac{t}{\delta^{L-2}}\right)\right)$
$=\mathcal{J}\left(\mathbf{X};\frac{1}{\delta}\mathbf{w}\left(\frac{t}{\delta^{L-2}}\right)\right)^{\top}\left(\mathbf{y}-\xi\left(\frac{t}{\delta^{L-2}}\right)\right)$
$=\mathcal{J}\left(\mathbf{X};\mathbf{s}(t)\right)^{\top}\left(\mathbf{y}-\xi\left(\frac{t}{\delta^{L-2}}\right)\right),$ (50)
where in third equality, from Lemma B.1, we used $(L-1)$-homogeneity of $\mathcal{J}\left(\mathbf{X};\mathbf{w}\right)$. Using $L$-homogeneity of $\mathcal{H}(\mathbf{X};\mathbf{w})$, we have, for square loss,
$\xi(t/\delta^{L-2})=\mathcal{H}(\mathbf{X};\mathbf{w}(t/\delta^{L-2}))=\delta^{L}\mathcal{H}(\mathbf{X};\mathbf{w}(t/\delta^{L-2})/\delta)=\delta^{L}\mathcal{H}(\mathbf{X};\mathbf{s}(t)),$
and similarly for logistic loss,
$\xi_{i}(t/\delta^{L-2})=y_{i}\left(\frac{1-e^{-y_{i}\delta^{L}\mathcal{H}(\mathbf{x}_{i};\mathbf{s}(t))}}{1+e^{-y_{i}\delta^{L}\mathcal{H}(\mathbf{x}_{i};\mathbf{s}(t))}}\right).$
Let $\tilde{\xi}(t)=\xi(t/\delta^{L-2})$, then
$\dot{\mathbf{s}}=\mathcal{J}\left(\mathbf{X};\mathbf{s}(t)\right)^{\top}(\mathbf{y}-\tilde{\xi}(t)),\mathbf{s}(0)=\mathbf{w}_{0}.$
Next, we will show that $\|\mathbf{s}(t)-\mathbf{u}(t)\|_{2}\leq\epsilon,$ for all $t\in[0,T_{\epsilon}]$. We first note that, $\mathbf{s}(0)=\mathbf{u}(0)=\mathbf{w}_{0}.$ Define
$\overline{T}_{\delta}=\inf_{t\in[0,T_{\epsilon}]}\{t:\|\mathbf{s}(t)-\mathbf{u}(t)\|_{2}=\epsilon\}.$
Here, $\overline{T}_{\delta}$ indicates the time when $\|\mathbf{s}(t)-\mathbf{u}(t)\|_{2}$ becomes $\epsilon$ for the first time, and $\delta$ in the subscript indicates that $\overline{T}_{\delta}$ could change with $\delta$. By definition of $\overline{T}_{\delta}$, for all $t\in[0,\overline{T}_{\delta}],$ $\|\mathbf{s}(t)-\mathbf{u}(t)\|_{2}\leq\epsilon$. Now, if we can show $\overline{T}_{\delta}=T_{\epsilon}$, then we would have proved $\|\mathbf{s}(t)-\mathbf{u}(t)\|_{2}\leq\epsilon,$ for all $t\in[0,T_{\epsilon}]$. Since we have assumed $\overline{T}_{\delta}\leq T_{\epsilon}$, we have that $\|\mathbf{u}(t)\|_{2}\leq B_{\epsilon}$ and $\|\mathbf{s}(t)\|_{2}\leq B_{\epsilon}+\epsilon=\tilde{B}_{\epsilon}$, for all $t\in[0,\overline{T}_{\delta}]$. Now, using (44), we have
$\left\|\mathcal{J}(\mathbf{X};\mathbf{u}(t))-\mathcal{J}(\mathbf{X};\mathbf{s}(t))\right\|_{2}\leq K_{\epsilon}\|\mathbf{u}(t)-\mathbf{s}(t)\|_{2},\forall t\in[0,\overline{T}_{\delta}].$ (51)
For square loss and $t\in[0,\overline{T}_{\delta}]$, we have
$\|\tilde{\xi}(t)\|_{2}=\delta^{L}\|\mathcal{H}(\mathbf{X};\mathbf{s}(t))\|_{2}\leq\delta^{L}\beta\|\mathbf{s}(t)\|_{2}^{L}\leq\delta^{L}\beta\tilde{B}_{\epsilon}^{L}.$
Since $\delta\leq 1/\tilde{B}_{\epsilon}$, then for logistic loss and $t\in[0,\overline{T}_{\delta}]$, we have
$|\tilde{\xi}_{i}(t)|=\left|y_{i}\left(\frac{1-e^{-y_{i}\delta^{L}\mathcal{H}(\mathbf{x}_{i};\mathbf{s}(t))}}{1+e^{-y_{i}\delta^{L}\mathcal{H}(\mathbf{x}_{i};\mathbf{s}(t))}}\right)\right|\leq|y_{i}|\left|1-e^{-y_{i}\delta^{L}\mathcal{H}(\mathbf{x}_{i};\mathbf{s}(t))}\right|\leq|y_{i}|e^{|y_{i}|\beta}|y_{i}\delta^{L}\mathcal{H}(\mathbf{x}_{i};\mathbf{s}(t))|\leq|y_{i}|^{2}e^{|y_{i}|\beta}\delta^{L}\beta\tilde{B}_{\epsilon}^{L},$
where the first inequality follows since $|1+e^{x}|\geq 1$. Next, since $\delta\leq 1/\tilde{B}_{\epsilon}$, we have that for all $t\in[0,\overline{T}_{\delta}]$, $\delta^{L}\mathcal{H}(\mathbf{x}_{i};\mathbf{s}(t))\leq\delta^{L}\beta\tilde{B}_{\epsilon}^{L}\leq\beta.$ We obtain the second inequality using the first-order Taylor’s approximation
$\left|1-e^{-y_{i}\delta^{L}\mathcal{H}(\mathbf{x}_{i};\mathbf{s}(t))}\right|\leq e^{|y_{i}|\beta}|y_{i}\delta^{L}\mathcal{H}(\mathbf{x}_{i};\mathbf{s}(t))|.$
For the last inequality we use $\delta^{L}\mathcal{H}(\mathbf{x}_{i};\mathbf{s}(t))\leq\delta^{L}\beta\tilde{B}_{\epsilon}^{L},$ for all $t\in[0,\overline{T}_{\delta}]$. Therefore, $\|\tilde{\xi}(t)\|_{2}\leq\|\mathbf{y}\|^{2}e^{\|\mathbf{y}\|\beta}\delta^{L}\beta\tilde{B}_{\epsilon}^{L}$, for all $t\in[0,\overline{T}_{\delta}].$ Hence, by the definition of $C_{\epsilon}$, for both square loss and logistic loss, we have
$\|\tilde{\xi}(t)\|_{2}\leq C_{\epsilon}\delta^{L},\text{ for all }t\in[0,\overline{T}_{\delta}].$ (52)
Thus, for any $t\in[0,\overline{T}_{\delta}]$
$\frac{1}{2}\frac{d\|\mathbf{s}-\mathbf{u}\|_{2}^{2}}{dt}$ $=(\mathbf{s}-\mathbf{u})^{\top}(\dot{\mathbf{s}}-\dot{\mathbf{u}})$
$=(\mathbf{s}-\mathbf{u})^{\top}(\mathcal{J}(\mathbf{X};\mathbf{s})-\mathcal{J}(\mathbf{X};\mathbf{u}))^{\top}\mathbf{y}-(\mathbf{s}-\mathbf{u})^{\top}\mathcal{J}(\mathbf{X};\mathbf{s})^{\top}\tilde{\xi}(t)$
$\leq K_{\epsilon}\|\mathbf{y}\|_{2}\|\mathbf{s}-\mathbf{u}\|_{2}^{2}+\alpha\|\mathbf{s}\|_{2}^{L-1}\|\mathbf{s}-\mathbf{u}\|_{2}\|\tilde{\xi}(t)\|_{2}$
$\leq K_{\epsilon}\|\mathbf{y}\|_{2}\|\mathbf{s}-\mathbf{u}\|_{2}^{2}+\alpha\epsilon\tilde{B}_{\epsilon}^{L-1}C_{\epsilon}\delta^{L},$
where the first inequality follows from (51) and (46). For the second inequality we used (52), and $\|\mathbf{s}(t)\|_{2}\leq\tilde{B}_{\epsilon}$, for all $t\in[0,\overline{T}_{\delta}]$. Integrating the above inequality on both sides from 00 to $t_{1}\in[0,\overline{T}_{\delta}]$, we have
[…]
Then, using the second Gronwall’s inequality in Lemma A.3, we have
$\|\mathbf{s}(\overline{T}_{\delta})-\mathbf{u}(\overline{T}_{\delta})\|_{2}^{2}\leq 2\alpha\epsilon\tilde{B}_{\epsilon}^{L-1}C_{\epsilon}\delta^{L}\overline{T}_{\delta}e^{2\overline{T}_{\delta}K_{\epsilon}\|\mathbf{y}\|_{2}}.$
For the sake of contradiction assume that $\overline{T}_{\delta}<T_{\epsilon}$. Then, by definition of $\overline{T}_{\delta}$, we know $\|\mathbf{s}(\overline{T}_{\delta})-\mathbf{u}(\overline{T}_{\delta})\|_{2}=\epsilon$. Using the above inequality, and since $\delta<\overline{\delta}$, where $\overline{\delta}$ satisfies (45), we have
$\epsilon^{2}=\|\mathbf{s}(\overline{T}_{\delta})-\mathbf{u}(\overline{T}_{\delta})\|_{2}^{2}\leq 2\alpha\epsilon\tilde{B}_{\epsilon}^{L-1}C_{\epsilon}\delta^{L}T_{\epsilon}e^{2T_{\epsilon}K_{\epsilon}\|\mathbf{y}\|_{2}}\leq\epsilon^{2}/2,$
which leads to a contradiction. Thus, $\overline{T}_{\delta}=T_{\epsilon}$. Since $\|\mathbf{s}(t)-\mathbf{u}(t)\|_{2}\leq\epsilon$, for all $t\in[0,T_{\epsilon}]$, we have
$\|\mathbf{w}(t/\delta^{L-2})\|_{2}\leq\delta(\|\mathbf{u}(t)\|_{2}+\epsilon)\leq\tilde{B}_{\epsilon}\delta,\forall t\in[0,T_{\epsilon}].$
We next prove the second part of the theorem. If $\mathbf{u}(t)$ converges to $\mathbf{0}$, then from (41) we know $\|\mathbf{u}(T_{\epsilon})\|_{2}\leq\epsilon$, which implies
$\left\|\mathbf{w}\left(\frac{T_{\epsilon}}{\delta^{L-2}}\right)\right\|_{2}=\delta\|\mathbf{s}(T_{\epsilon})\|_{2}\leq 2\delta\epsilon.$ (53)
Else, from (42) and (43), we know
$\frac{\mathbf{u}(T_{\epsilon})^{\top}{\mathbf{u}^{*}}}{\|\mathbf{u}(T_{\epsilon})\|_{2}}\geq 1-\epsilon,\text{ and }\|\mathbf{u}(T_{\epsilon})\|_{2}\geq\eta.$ (54)
Define $\overline{T}_{\epsilon}=T_{\epsilon}/\delta^{L-2}$. Since $\left\|\frac{1}{\delta}{\mathbf{w}\left(\frac{T_{\epsilon}}{\delta^{L-2}}\right)}-\mathbf{u}(T_{\epsilon})\right\|_{2}\leq\epsilon$, then for some $\zeta\in\mathbb{R}^{k}$ we have
$\frac{\mathbf{w}(\overline{T}_{\epsilon})}{\delta}=\mathbf{u}(T_{\epsilon})+\zeta,$ (55)
where $\|\zeta\|_{2}\leq\epsilon$. From $\epsilon\in(0,\eta/2)$ and $\|\mathbf{u}(T_{\epsilon})\|_{2}\geq\eta$, we have $\|\mathbf{u}(T_{\epsilon})+\zeta\|_{2}\geq\eta/2$, which implies
$\left\|\mathbf{w}\left(\overline{T}_{\epsilon}\right)\right\|_{2}\geq\delta\eta/2.$
Next, from (55) we have
$\frac{\mathbf{w}(\overline{T}_{\epsilon})}{\|\mathbf{w}(\overline{T}_{\epsilon})\|_{2}}=\frac{\mathbf{u}(T_{\epsilon})+\zeta}{\|\mathbf{u}(T_{\epsilon})+\zeta\|_{2}}.$
Hence,
$\frac{\mathbf{w}(\overline{T}_{\epsilon})^{\top}{\mathbf{u}}^{*}}{\|\mathbf{w}(\overline{T}_{\epsilon})\|_{2}}$ $=\frac{\mathbf{u}(T_{\epsilon})^{\top}{\mathbf{u}}^{*}+\zeta^{\top}{\mathbf{u}}^{*}}{\|\mathbf{u}(T_{\epsilon})+\zeta\|_{2}}$
$=\left(\frac{\mathbf{u}(T_{\epsilon})^{\top}{\mathbf{u}}^{*}}{\|\mathbf{u}(T_{\epsilon})\|_{2}}\right)\frac{\|\mathbf{u}(T_{\epsilon})\|_{2}}{\|\mathbf{u}(T_{\epsilon})+\zeta\|_{2}}+\frac{\zeta^{\top}{\mathbf{u}}^{*}}{\|\mathbf{u}(T_{\epsilon})+\zeta\|_{2}}$
$\geq(1-\epsilon)\frac{\|\mathbf{u}(T_{\epsilon})\|_{2}}{\|\mathbf{u}(T_{\epsilon})\|_{2}+\|\zeta\|_{2}}-\frac{2\epsilon}{\eta}$
$=(1-\epsilon)\frac{1}{1+\frac{\|\zeta\|_{2}}{\|\mathbf{u}(T_{\epsilon})\|_{2}}}-\frac{2\epsilon}{\eta}$
$\geq\frac{1-\epsilon}{1+\frac{\epsilon}{\eta}}-\frac{2\epsilon}{\eta}\geq(1-\epsilon)\left(1-\frac{\epsilon}{\eta}\right)-\frac{2\epsilon}{\eta}\geq 1-\left(1+\frac{3}{\eta}\right)\epsilon,$
where in the first inequality we used (54) along with the fact that $\|\zeta\|_{2}\leq\epsilon$ and $\|\mathbf{u}(T_{\epsilon})+\zeta\|_{2}\geq\eta/2.$ The second inequality follows since $\|\zeta\|_{2}\leq\epsilon$ and $\|\mathbf{u}(T_{\epsilon})\|_{2}\geq\eta.$
Appendix C Proofs Omitted from Section 5.2
C.1 Proof of Theorem 5.3
Proof.
Since $\{\overline{\mathbf{w}}_{n},\overline{\mathbf{w}}_{z}\}$ is a saddle point of
$\mathcal{L}\left(\mathbf{w}_{n},\mathbf{w}_{z}\right)=\frac{1}{2}\left\|\mathcal{H}_{n}(\mathbf{X};\mathbf{w}_{n})+\mathcal{H}_{z}(\mathbf{X};\mathbf{w}_{z})-\mathbf{y}\right\|^{2},$ (56)
it follows that
$\begin{bmatrix}\mathbf{0}\\ \mathbf{0}\end{bmatrix}=\begin{bmatrix}\nabla\mathcal{H}_{n}(\mathbf{X};\overline{\mathbf{w}}_{n})^{\top}\overline{\mathbf{y}}\\ \nabla\mathcal{H}_{z}(\mathbf{X};\overline{\mathbf{w}}_{z})^{\top}\overline{\mathbf{y}},\end{bmatrix},$
where $\overline{\mathbf{y}}=\mathbf{y}-\mathcal{H}_{n}(\mathbf{X};\overline{\mathbf{w}}_{n})$. Also, we know that if $\|\mathbf{w}_{n}-\overline{\mathbf{w}}_{n}\|_{2}\leq\gamma,$ then
$\left\langle\mathbf{w}_{n}-\overline{\mathbf{w}}_{n},\nabla_{\mathbf{w}_{n}}\mathcal{L}(\mathbf{w}_{n},\mathbf{0})\right\rangle\geq 0.$
In this section, $\mathcal{J}_{z}\left(\mathbf{X};\mathbf{w}_{z}\right)$ and $\mathcal{J}_{n}\left(\mathbf{X};\mathbf{w}_{n}\right)$ denotes the Jacobian of $\mathcal{H}_{z}(\mathbf{X};\mathbf{w}_{z})$ and $\mathcal{H}_{n}(\mathbf{X};\mathbf{w}_{n})$, respectively. Then, the above inequality can be written as
$\left\langle\mathbf{w}_{n}-\overline{\mathbf{w}}_{n},\mathcal{J}_{n}\left(\mathbf{X};\mathbf{w}_{n}\right)^{\top}(\mathcal{H}_{n}(\mathbf{X};\mathbf{w}_{n})-\mathbf{y})\right\rangle\geq 0.$ (57)
Let $\mathbf{p}_{z}$ be a fixed unit norm vector and $\mathbf{u}(t)$ be the solution of
$\frac{d\mathbf{u}}{dt}=\nabla\mathcal{N}_{\overline{\mathbf{y}},\mathcal{H}_{z}}(\mathbf{u})=\mathcal{J}_{z}\left(\mathbf{X};\mathbf{u}\right)^{\top}\overline{\mathbf{y}},\mathbf{u}(0)=\mathbf{p}_{z}.$
Similar to the proof of Theorem 5.1, we next define $\eta$, $T_{\epsilon}$ and $B_{\epsilon}$. From Lemma 5.2, we know $\mathbf{u}(t)$ either converges to $\mathbf{0}$, or $\mathbf{u}(t)$ converges in direction to a non-negative KKT point of constrained NCF defined with respect to $\overline{\mathbf{y}}$ and $\mathcal{H}_{z}$. If $\mathbf{u}(t)$ converges to $\mathbf{0}$, then we define $\eta=\min(2,2\gamma).$ Then, for any $\epsilon\in(0,\eta/2)$, we define $B_{\epsilon}=\epsilon$ and choose $T_{\epsilon}$ large enough such that
$\|\mathbf{u}(T_{\epsilon})\|_{2}\leq B_{\epsilon}=\epsilon.$ (58)
Else, if $\mathbf{u}(t)$ does not converge to $\mathbf{0}$, then from Lemma B.6, there exists $\tilde{\eta}>0$ and $T$ such that $\|\mathbf{u}(t)\|_{2}\geq\tilde{\eta}$, for all $t\geq T$. We next define $\eta=\min(\tilde{\eta},1,2\gamma)$. Then, from Lemma B.6, for any $\epsilon\in(0,\eta/2)$ there exists $T_{\epsilon}\geq T$ and $B_{\epsilon}$ such that
$\frac{\mathbf{u}(T_{\epsilon})^{\top}{\mathbf{u}^{*}}}{\|\mathbf{u}(T_{\epsilon})\|_{2}}\geq 1-\epsilon,\text{ and }\|\mathbf{u}(T_{\epsilon})\|_{2}\leq B_{\epsilon},\forall t\leq T_{\epsilon},$ (59)
where $\mathbf{u}^{*}$ is a non-negative KKT point of
$\max_{\|\mathbf{u}\|_{2}^{2}=1}\mathcal{N}_{\overline{\mathbf{y}},\mathcal{H}_{z}}(\mathbf{u}).$
Additionally, since $T_{\epsilon}\geq T$, we have
$\|\mathbf{u}(T_{\epsilon})\|_{2}\geq\tilde{\eta}\geq\eta.$ (60)
We define
$\beta=\max\left(\sup\{\|\mathcal{H}_{n}(\mathbf{X};\mathbf{w}_{n})\|_{2}:\|\mathbf{w}_{n}\|_{2}=1\},\sup\{\|\mathcal{H}_{z}(\mathbf{X};\mathbf{w}_{z})\|_{2}:\|\mathbf{w}_{z}\|_{2}=1\}\right).$ (61)
and
$\alpha=\max\left(\sup\{\|\mathcal{J}_{n}(\mathbf{X};\mathbf{w}_{n})\|_{2}:\|\mathbf{w}_{n}\|_{2}=1\},\sup\{\|\mathcal{J}_{z}(\mathbf{X};\mathbf{w}_{z})\|_{2}:\|\mathbf{w}_{z}\|_{2}=1\}\right).$ (62)
Next, we define other $\epsilon$ dependent parameters which are useful in our proof. Let $\tilde{B}_{\epsilon}=B_{\epsilon}+\epsilon,\tilde{W}_{\epsilon}=\|\overline{\mathbf{w}}_{n}\|_{2}+\epsilon$, and $C_{\epsilon}=(1+2\epsilon\alpha\tilde{W}_{\epsilon}^{L-1}\beta\tilde{B}_{\epsilon}^{L}T_{\epsilon})$. Since $\mathcal{H}_{z}(\mathbf{x};\mathbf{w}_{z})$ has locally Lipschitz gradient, there exists $K_{\epsilon}^{1}$ such that if $\|\mathbf{w}_{1z}\|_{2},\|\mathbf{w}_{2z}\|_{2}\leq\tilde{B}_{\epsilon}$, then
$\left\|\mathcal{J}_{z}(\mathbf{X};\mathbf{w}_{1z})-\mathcal{J}_{z}(\mathbf{X};\mathbf{w}_{2z})\right\|_{2}\leq K^{1}_{\epsilon}\|\mathbf{w}_{1z}-\mathbf{w}_{2z}\|_{2}.$ (63)
Similarly, by local Lipschitz property of $\mathcal{H}_{n}(\mathbf{X};\mathbf{w}_{n})$ there exists $K_{\epsilon}^{2}$ such that if $\|\mathbf{w}_{1n}\|_{2}\leq\tilde{W}_{\epsilon}$, then
$\left\|\mathcal{H}_{n}(\mathbf{X};\mathbf{w}_{1n})-\mathcal{H}_{n}(\mathbf{X};\overline{\mathbf{w}}_{n})\right\|_{2}\leq K^{2}_{\epsilon}\|\mathbf{w}_{1n}-\overline{\mathbf{w}}_{n}\|_{2}.$ (64)
We further define
$\bar{\delta}^{2}=\min\left(1,\epsilon^{2},\frac{\epsilon^{2}}{2C_{\epsilon}},\frac{\epsilon}{4\alpha\tilde{B}_{\epsilon}^{L-1}(K_{\epsilon}^{2}C_{\epsilon}+\beta\tilde{B}_{\epsilon}^{L})T_{\epsilon}e^{2T_{\epsilon}K^{1}_{\epsilon}\|\overline{\mathbf{y}}\|_{2}}}\right).$ (65)
Now, for any $\delta\in(0,\overline{\delta})$, let $\{\mathbf{w}_{n}(t),\mathbf{w}_{z}(t)\}$ satisfy for all $t\geq 0$
$\begin{bmatrix}\dot{\mathbf{w}}_{n}\\ \dot{\mathbf{w}}_{z}\end{bmatrix}=\begin{bmatrix}\mathcal{J}_{n}(\mathbf{X};\mathbf{w}_{n})^{\top}(\mathbf{y}-\mathcal{H}_{n}(\mathbf{X};\mathbf{w}_{n})+\mathcal{H}_{z}(\mathbf{X};\mathbf{w}_{z}))\\ \mathcal{J}_{z}(\mathbf{X};\mathbf{w}_{z})^{\top}(\mathbf{y}-\mathcal{H}_{n}(\mathbf{X};\mathbf{w}_{n})+\mathcal{H}_{z}(\mathbf{X};\mathbf{w}_{z}))\end{bmatrix},\begin{bmatrix}{\mathbf{w}}_{n}(0)\\ {\mathbf{w}_{z}}(0)\end{bmatrix}=\begin{bmatrix}\overline{\mathbf{w}}_{n}+\delta\mathbf{p}_{n}\\ \overline{\mathbf{w}}_{z}+\delta\mathbf{p}_{z}\end{bmatrix},$ (74)
where $\mathbf{p}_{n}$ and $\mathbf{p}_{z}$ are unit norm vectors. We define
$\mathbf{s}_{n}(t)=\mathbf{w}_{n}\left(\frac{t}{\delta^{L-2}}\right)\text{ and }\mathbf{s}_{z}(t)=\frac{1}{\delta}\mathbf{w}_{z}\left(\frac{t}{\delta^{L-2}}\right),$
where note that in contrast to $\mathbf{s}_{z}(t)$, $\mathbf{s}_{n}(t)$ does not have a factor of $1/\delta$. Then,
$\frac{d\mathbf{s}_{n}}{dt}$ $=\frac{1}{\delta^{L-2}}\dot{\mathbf{w}}_{n}\left(\frac{t}{\delta^{L-2}}\right)$
$=\frac{1}{\delta^{L-2}}\mathcal{J}_{n}\left(\mathbf{X};\mathbf{w}_{n}\left(\frac{t}{\delta^{L-2}}\right)\right)^{\top}\left(\mathbf{y}-\mathcal{H}_{n}\left(\mathbf{X};\mathbf{w}_{n}\left(\frac{t}{\delta^{L-2}}\right)\right)-\mathcal{H}_{z}\left(\mathbf{X};\mathbf{w}_{z}\left(\frac{t}{\delta^{L-2}}\right)\right)\right)$
$=\frac{1}{\delta^{L-2}}\mathcal{J}_{n}\left(\mathbf{X};\mathbf{s}_{n}\left(t\right)\right)^{\top}(\mathbf{y}-\mathcal{H}_{n}\left(\mathbf{X};\mathbf{s}_{n}\left(t\right)\right)-\delta^{L}\mathcal{H}_{z}\left(\mathbf{X};\mathbf{s}_{z}\left(t\right)\right)),$ (75)
and
$\frac{d\mathbf{s}_{z}}{dt}$ $=\frac{1}{\delta^{L-1}}\dot{\mathbf{w}}_{z}\left(\frac{t}{\delta^{L-2}}\right)$
$=\frac{1}{\delta^{L-1}}\mathcal{J}_{z}\left(\mathbf{X};\mathbf{w}_{z}\left(\frac{t}{\delta^{L-2}}\right)\right)^{\top}\left(\mathbf{y}-\mathcal{H}_{n}\left(\mathbf{X};\mathbf{w}_{n}\left(\frac{t}{\delta^{L-2}}\right)\right)-\mathcal{H}_{z}\left(\mathbf{X};\mathbf{w}_{z}\left(\frac{t}{\delta^{L-2}}\right)\right)\right)$
$=\mathcal{J}_{z}\left(\mathbf{X};\mathbf{s}_{z}\left(t\right)\right)^{\top}(\mathbf{y}-\mathcal{H}_{n}\left(\mathbf{X};\mathbf{s}_{n}\left(t\right)\right)-\delta^{L}\mathcal{H}_{z}\left(\mathbf{X};\mathbf{s}_{z}\left(t\right)\right)),$ (76)
where in the above set of equalities we mainly used $(L-1)$-homogeneity of the Jacobians and $L$-homogeneity of the neural network. We next define
$\mathbf{Z}(t)=\max(\|\mathbf{s}_{n}(t)-\overline{\mathbf{w}}_{n}\|_{2}^{2},\|\mathbf{s}_{z}(t)-\mathbf{u}(t)\|_{2}^{2}).$
Next, we will show that $\mathbf{Z}(t)\leq\epsilon$, for all $t\in[0,T_{\epsilon}]$. Note that $\mathbf{Z}(0)=\delta^{2}<\epsilon^{2}\leq\gamma^{2}$. Define
$\overline{T}_{\delta}=\inf_{t\in[0,T_{\epsilon}]}\{t:\mathbf{Z}(t)=\epsilon^{2}\}.$
Here, $\overline{T}_{\delta}$ indicates the time when $\mathbf{Z}(t)$ becomes $\epsilon^{2}$ for the first time. Thus, for all $t\in[0,\overline{T}_{\delta}],$ $\mathbf{Z}(t)\leq\epsilon^{2}\leq\gamma^{2}$. Note that if we can show $\overline{T}_{\delta}=T_{\epsilon}$, then it would imply $\mathbf{Z}(t)\leq\epsilon$, for all $t\in[0,T_{\epsilon}]$. Since $\overline{T}_{\delta}\leq T_{\epsilon}$, we have that $\|\mathbf{u}(t)\|_{2}\leq B_{\epsilon}$ and $\|\mathbf{s}_{z}(t)\|_{2}\leq B_{\epsilon}+\epsilon=\tilde{B}_{\epsilon}$, for all $t\in[0,\overline{T}_{\delta}]$. Also, $\|\mathbf{s}_{n}(t)\|_{2}\leq\|\overline{\mathbf{w}}_{n}\|_{2}+\epsilon=\tilde{W}_{\epsilon}$. From (63) and (64), for all $t\in[0,\overline{T}_{\delta}]$, we have
$\left\|\mathcal{J}_{z}(\mathbf{X};\mathbf{u}(t))-\mathcal{J}_{z}(\mathbf{X};\mathbf{s}_{z}(t))\right\|_{2}\leq K^{1}_{\epsilon}\|\mathbf{u}(t)-\mathbf{s}_{z}(t)\|_{2},$ (77)
and
$\left\|\mathcal{H}_{n}(\mathbf{X};\mathbf{s}_{n}(t))-\mathcal{H}_{n}(\mathbf{X};\overline{\mathbf{w}}_{n})\right\|_{2}\leq K^{2}_{\epsilon}\|\mathbf{s}_{n}(t)-\overline{\mathbf{w}}_{n}\|_{2}.$ (78)
Next, for all $t\in[0,\overline{T}_{\delta}]$, we have
$\frac{1}{2}\frac{d\left\|{\mathbf{s}}_{n}-\overline{\mathbf{w}}_{n}\right\|^{2}_{2}}{dt}$ $=(\mathbf{s}_{n}-\overline{\mathbf{w}}_{n})^{\top}\dot{\mathbf{s}}_{n}$
$=\frac{1}{\delta^{L-2}}(\mathbf{s}_{n}-\overline{\mathbf{w}}_{n})^{\top}\mathcal{J}_{n}\left(\mathbf{X};\mathbf{s}_{n}\right)^{\top}(\mathbf{y}-\mathcal{H}_{n}\left(\mathbf{X};\mathbf{s}_{n}\right)-\delta^{L}\mathcal{H}_{z}\left(\mathbf{X};\mathbf{s}_{z}\right))$
$\leq\frac{1}{\delta^{L-2}}(\mathbf{s}_{n}-\overline{\mathbf{w}}_{n})^{\top}\mathcal{J}_{n}\left(\mathbf{X};\mathbf{s}_{n}\right)^{\top}(-\delta^{L}\mathcal{H}_{z}\left(\mathbf{X};\mathbf{s}_{z}\right))$
$\leq\delta^{2}\|\mathbf{s}_{n}-\overline{\mathbf{w}}_{n}\|_{2}\|\mathcal{J}_{n}\left(\mathbf{X};\mathbf{s}_{n}\right)\|_{2}\|\mathcal{H}_{z}\left(\mathbf{X};\mathbf{s}_{z}\right)\|_{2}$
$\leq\delta^{2}\epsilon\alpha\|\mathbf{s}_{n}\|_{2}^{L-1}\beta\|\mathbf{s}_{z}\|_{2}^{L}\leq\delta^{2}\epsilon\alpha\tilde{W}_{\epsilon}^{L-1}\beta\tilde{B}_{\epsilon}^{L},$
where the first inequality follows from (57) and that $\left\|{\mathbf{s}}_{n}(t)-\overline{\mathbf{w}}_{n}\right\|_{2}\leq\epsilon\leq\gamma$, for all $t\in[0,\overline{T}_{\delta}]$. The subsequent set of inequalities follow by Cauchy-Schwartz, and using $\|\mathbf{s}_{z}(t)\|_{2}\leq\tilde{B}_{\epsilon}$ and $\|\mathbf{s}_{n}(t)\|_{2}\leq\tilde{W}_{\epsilon}$, for all $t\in[0,\overline{T}_{\delta}].$ Since $\overline{T}_{\delta}\leq T_{\epsilon}$, integrating the above inequality from 00 to $t\in[0,\overline{T}_{\delta}]$ gives us
$\left\|{\mathbf{s}}_{n}(t)-\overline{\mathbf{w}}_{n}\right\|^{2}_{2}\leq\left\|{\mathbf{s}}_{n}(0)-\overline{\mathbf{w}}_{n}\right\|^{2}_{2}+2\delta^{2}\epsilon\alpha\tilde{W}_{\epsilon}^{L-1}\beta\tilde{B}_{\epsilon}^{L}T_{\epsilon}\leq\delta^{2}(1+2\epsilon\alpha\tilde{W}_{\epsilon}^{L-1}\beta\tilde{B}_{\epsilon}^{L}T_{\epsilon})=C_{\epsilon}\delta^{2}.$ (79)
Next, for all $t\in[0,\overline{T}_{\delta}]$, we have
$\frac{1}{2}\frac{d\|\mathbf{s}_{z}-\mathbf{u}\|_{2}^{2}}{dt}=(\mathbf{s}_{z}-\mathbf{u})^{\top}(\dot{\mathbf{s}}_{z}-\dot{\mathbf{u}})$
$=$ $(\mathbf{s}_{z}-\mathbf{u})^{\top}(\mathcal{J}_{z}(\mathbf{X};\mathbf{s}_{z})^{\top}(\mathbf{y}-\mathcal{H}_{n}(\mathbf{X};\mathbf{s}_{n})-\delta^{L}\mathcal{H}_{z}(\mathbf{X};\mathbf{s}_{z}))-\mathcal{J}_{z}(\mathbf{X};\mathbf{u})^{\top}\overline{\mathbf{y}})$
$=$ $(\mathbf{s}_{z}-\mathbf{u})^{\top}(\mathcal{J}_{z}(\mathbf{X};\mathbf{s}_{z})-\mathcal{J}_{z}(\mathbf{X};\mathbf{u}))^{\top}\overline{\mathbf{y}}+(\mathbf{s}_{z}-\mathbf{u})^{\top}\mathcal{J}_{z}(\mathbf{X};\mathbf{s}_{z})^{\top}\left(\mathcal{H}_{n}(\mathbf{X};\overline{\mathbf{w}}_{n})-\mathcal{H}_{n}(\mathbf{X};\mathbf{s}_{n})-\delta^{L}\mathcal{H}_{z}(\mathbf{X};\mathbf{s}_{z})\right)$
$\leq$ $K^{1}_{\epsilon}\|\overline{\mathbf{y}}\|_{2}\|\mathbf{s}_{z}-\mathbf{u}\|_{2}^{2}+\alpha\|\mathbf{s}_{z}(t)\|_{2}^{L-1}\|\mathbf{s}_{z}-\mathbf{u}\|_{2}\left(\|\mathcal{H}_{n}(\mathbf{X};\overline{\mathbf{w}}_{n})-\mathcal{H}_{n}(\mathbf{X};\mathbf{s}_{n})\|_{2}+\beta\delta^{L}\|\mathbf{s}_{z}\|_{2}^{L}\right)$
$\leq$ $K^{1}_{\epsilon}\|\overline{\mathbf{y}}\|_{2}\|\mathbf{s}_{z}-\mathbf{u}\|_{2}^{2}+\alpha\|\mathbf{s}_{z}(t)\|_{2}^{L-1}\|\mathbf{s}_{z}-\mathbf{u}\|_{2}\left(K_{\epsilon}^{2}C_{\epsilon}\delta^{2}+\beta\tilde{B}_{\epsilon}^{L}\delta^{L}\right)$
$\leq$ $K^{1}_{\epsilon}\|\overline{\mathbf{y}}\|_{2}\|\mathbf{s}_{z}-\mathbf{u}\|_{2}^{2}+\alpha\epsilon\tilde{B}_{\epsilon}^{L-1}(K_{\epsilon}^{2}C_{\epsilon}\delta^{2}+\beta\tilde{B}_{\epsilon}^{L}\delta^{L})$
$\leq$ $K^{1}_{\epsilon}\|\overline{\mathbf{y}}\|_{2}\|\mathbf{s}_{z}-\mathbf{u}\|_{2}^{2}+\alpha\epsilon\tilde{B}_{\epsilon}^{L-1}(K_{\epsilon}^{2}C_{\epsilon}\delta^{2}+\beta\tilde{B}_{\epsilon}^{L}\delta^{2}),$
where in the last equality we used $\overline{\mathbf{y}}=\mathbf{y}-\mathcal{H}_{n}(\mathbf{X};\overline{\mathbf{w}}_{n}).$ In the first inequality we used (77) and the Cauchy-Schwartz inequality. The second inequality follows from (78) and (79), and since $\|\mathbf{s}_{z}(t)\|_{2}\leq\tilde{B}_{\epsilon}$, for all $t\in[0,\overline{T}_{\delta}].$ In last inequality we used $\delta\leq 1.$ Then, integrating the above inequality and using the second Gronwall’s inequality in Lemma A.3, we have
$\|\mathbf{s}_{z}(\overline{T}_{\delta})-\mathbf{u}(\overline{T}_{\delta})\|_{2}^{2}\leq 2\alpha\epsilon\tilde{B}_{\epsilon}^{L-1}(K_{\epsilon}^{2}C_{\epsilon}+\beta\tilde{B}_{\epsilon}^{L})\delta^{2}\overline{T}_{\delta}e^{2\overline{T}_{\delta}K^{1}_{\epsilon}\|\overline{\mathbf{y}}\|_{2}}.$ (80)
For the sake of contradiction assume that $\overline{T}_{\delta}<T_{\epsilon}$. Then, by definition of $\overline{T}_{\delta}$, atleast one of the following equalities must hold
$\|\mathbf{s}_{z}(\overline{T}_{\delta})-\mathbf{u}(\overline{T}_{\delta})\|_{2}^{2}=\epsilon^{2}\mbox{ or }\left\|{\mathbf{s}}_{n}(\overline{T}_{\delta})-\overline{\mathbf{w}}_{n}\right\|^{2}_{2}=\epsilon^{2}$
However, from (79), and since $\delta<\overline{\delta}$, where $\overline{\delta}$ satisfies (65), we have
$\left\|{\mathbf{s}}_{n}(\overline{T}_{\delta})-\overline{\mathbf{w}}_{n}\right\|^{2}_{2}\leq C_{\epsilon}\delta^{2}\leq\epsilon^{2}/2,$
and from (80), we have
$\|\mathbf{s}_{z}(\overline{T}_{\delta})-\mathbf{u}(\overline{T}_{\delta})\|_{2}^{2}$ $\leq 2\alpha\epsilon\tilde{B}_{\epsilon}^{L-1}(K_{\epsilon}^{2}C_{\epsilon}+\beta\tilde{B}_{\epsilon}^{L})\delta^{2}\overline{T}_{\delta}e^{2\overline{T}_{\delta}K^{1}_{\epsilon}\|\overline{\mathbf{y}}\|_{2}}$
$\leq 2\alpha\epsilon\tilde{B}_{\epsilon}^{L-1}(K_{\epsilon}^{2}C_{\epsilon}+\beta\tilde{B}_{\epsilon}^{L})\delta^{2}T_{\epsilon}e^{2T_{\epsilon}K^{1}_{\epsilon}\|\overline{\mathbf{y}}\|_{2}}$
$\leq\epsilon^{2}/2,$
which leads to a contradiction. Thus, $\overline{T}_{\delta}=T_{\epsilon}$. Define $\tilde{C}_{\epsilon}=\sqrt{C_{\epsilon}}$, then from (79), for all $t\in[0,T_{\epsilon}]$, we have
$\left\|{\mathbf{s}}_{n}(t)-\overline{\mathbf{w}}_{n}\right\|_{2}\leq\tilde{C}_{\epsilon}\delta,$
and
$\|\mathbf{s}_{z}(t)-\mathbf{u}(t)\|_{2}\leq\epsilon\implies\|\mathbf{w}_{z}(t/\delta^{L-2})\|_{2}\leq\tilde{B}_{\epsilon}\delta,$ (81)
which proves the first part of the theorem. We next prove the second part of the theorem. Define $\overline{T}_{\epsilon}=T_{\epsilon}/\delta^{L-2}$, then from (81), we have $\|\mathbf{w}_{z}(\overline{T}_{\epsilon})/\delta-\mathbf{u}(T_{\epsilon})\|_{2}\leq\epsilon$. Thus, if $\mathbf{u}(t)$ converges to $\mathbf{0}$, then, from (58) we have $\|\mathbf{u}(T_{\epsilon})\|_{2}\leq\epsilon$, which implies
$\left\|\mathbf{w}_{z}\left(\overline{T}_{\epsilon}\right)\right\|_{2}\leq 2\delta\epsilon.$ (82)
Else, from (59) and (60), we have
$\frac{\mathbf{u}(T_{\epsilon})^{\top}{\mathbf{u}^{*}}}{\|\mathbf{u}(T_{\epsilon})\|_{2}}\geq 1-\epsilon,\text{ and }\|\mathbf{u}(T_{\epsilon})\|_{2}\geq\eta.$ (83)
Since $\left\|\frac{1}{\delta}{\mathbf{w}_{z}\left(\overline{T}_{\epsilon}\right)}-\mathbf{u}(T_{\epsilon})\right\|_{2}\leq\epsilon$, for some vector $\zeta$ we have
$\frac{\mathbf{w}_{z}(\overline{T}_{\epsilon})}{\delta}=\mathbf{u}(T_{\epsilon})+\zeta,$ (84)
where $\|\zeta\|_{2}\leq\epsilon$. From $\epsilon\in(0,\eta/2)$ and $\|\mathbf{u}(T_{\epsilon})\|_{2}\geq\tilde{\eta}\geq\eta$, we have $\|\mathbf{u}(T_{\epsilon})+\zeta\|_{2}\geq\eta/2$, which implies
$\|\mathbf{w}_{z}(\overline{T}_{\epsilon})\|_{2}\geq\delta\eta/2.$
Next, from (84) we have
$\frac{\mathbf{w}_{z}(\overline{T}_{\epsilon})}{\|\mathbf{w}_{z}(\overline{T}_{\epsilon})\|_{2}}=\frac{\mathbf{u}(T_{\epsilon})+\zeta}{\|\mathbf{u}(T_{\epsilon})+\zeta\|_{2}}.$
Hence,
$\frac{\mathbf{w}_{z}(\overline{T}_{\epsilon})^{\top}{\mathbf{u}}^{*}}{\|\mathbf{w}_{z}(\overline{T}_{\epsilon})\|_{2}}$ $=\frac{\mathbf{u}(T_{\epsilon})^{\top}{\mathbf{u}}^{*}+\zeta^{\top}{\mathbf{u}}^{*}}{\|\mathbf{u}(T_{\epsilon})+\zeta\|_{2}}$
$=\left(\frac{\mathbf{u}(T_{\epsilon})^{\top}{\mathbf{u}}^{*}}{\|\mathbf{u}(T_{\epsilon})\|_{2}}\right)\frac{\|\mathbf{u}(T_{\epsilon})\|_{2}}{\|\mathbf{u}(T_{\epsilon})+\zeta\|_{2}}+\frac{\zeta^{\top}{\mathbf{u}}^{*}}{\|\mathbf{u}(T_{\epsilon})+\zeta\|_{2}}$
$\geq(1-\epsilon)\frac{\|\mathbf{u}(T_{\epsilon})\|_{2}}{\|\mathbf{u}(T_{\epsilon})\|_{2}+\|\zeta\|_{2}}-\frac{2\epsilon}{\eta}$
$=(1-\epsilon)\frac{1}{1+\frac{\|\zeta\|_{2}}{\|\mathbf{u}(T_{\epsilon})\|_{2}}}-\frac{2\epsilon}{\eta}$
$\geq\frac{1-\epsilon}{1+\frac{\epsilon}{\eta}}-\frac{2\epsilon}{\eta}\geq(1-\epsilon)\left(1-\frac{\epsilon}{\eta}\right)-\frac{2\epsilon}{\eta}\geq 1-\left(1+\frac{3}{\eta}\right)\epsilon,$
where in the first inequality we used (83) along with the fact that $\|\zeta\|_{2}\leq\epsilon$ and $\|\mathbf{u}(T_{\epsilon})+\zeta\|_{2}\geq\eta/2.$ The second inequality follows since $\|\zeta\|_{2}\leq\epsilon$ and $\|\mathbf{u}(T_{\epsilon})\|_{2}\geq\eta.$
C.2 Proof of Lemma 5.4
Proof.
We begin by proving that $\overline{\mathbf{w}}_{n}$ is a saddle point of
$\mathcal{L}\left(\mathbf{w}_{n},\mathbf{0}\right)=\frac{1}{2}\left\|\mathcal{H}_{n}(\mathbf{X};\mathbf{w}_{n})-\mathbf{y}\right\|^{2}.$ (85)
Note that $\nabla_{\mathbf{w}_{n}}\mathcal{L}\left({\mathbf{w}}_{n},\mathbf{0}\right)=\mathcal{J}_{n}(\mathbf{X};{\mathbf{w}}_{n})^{\top}\left(\mathcal{H}_{n}(\mathbf{X};{\mathbf{w}}_{n})-\mathbf{y}\right).$ Since $\{\overline{\mathbf{w}}_{n},\overline{\mathbf{w}}_{z}\}$ is a saddle point of
$\mathcal{L}\left(\mathbf{w}_{n},\mathbf{w}_{z}\right)=\frac{1}{2}\left\|\mathcal{H}_{n}(\mathbf{X};\mathbf{w}_{n})+\mathcal{H}_{z}(\mathbf{X};\mathbf{w}_{z})-\mathbf{y}\right\|^{2},$ (86)
we have
$\begin{bmatrix}\mathbf{0}\\ \mathbf{0}\end{bmatrix}=\begin{bmatrix}\mathcal{J}_{n}(\mathbf{X};\overline{\mathbf{w}}_{n})^{\top}\overline{\mathbf{y}}\\ \mathcal{J}_{z}(\mathbf{X};\overline{\mathbf{w}}_{z})^{\top}\overline{\mathbf{y}}\end{bmatrix},$ (91)
where $\overline{\mathbf{y}}=\mathbf{y}-\mathcal{H}_{n}(\mathbf{X};\overline{\mathbf{w}}_{n})+\mathcal{H}_{z}(\mathbf{X};\overline{\mathbf{w}}_{z})$. Since $\overline{\mathbf{w}}_{z}=\mathbf{0}$ implies $\mathcal{H}_{z}(\mathbf{X};\overline{\mathbf{w}}_{z})=0$, we have
$\mathcal{J}_{n}(\mathbf{X};\overline{\mathbf{w}}_{n})^{\top}\left(\mathbf{y}-\mathcal{H}_{n}(\mathbf{X};\overline{\mathbf{w}}_{n})\right)=\mathbf{0}\implies\nabla_{\mathbf{w}_{n}}\mathcal{L}\left(\overline{\mathbf{w}}_{n},\mathbf{0}\right)=\mathbf{0},$
proving our first result. Next, suppose there exists $\gamma>0$ such that
$\left\langle\mathbf{w}_{n}-\overline{\mathbf{w}}_{n},\nabla_{\mathbf{w}_{n}}\mathcal{L}\left({\mathbf{w}}_{n},\mathbf{0}\right)\right\rangle\geq 0,\text{ if }\|\mathbf{w}_{n}-\overline{\mathbf{w}}_{n}\|_{2}\leq\gamma.$ (92)
Choose any $\mathbf{w}_{n}$ such that $\|\mathbf{w}_{n}-\overline{\mathbf{w}}_{n}\|_{2}\leq\gamma$. By the mean value theorem we have
$\mathcal{L}\left(\overline{\mathbf{w}}_{n},\mathbf{0}\right)=\mathcal{L}\left({\mathbf{w}}_{n},\mathbf{0}\right)+\left\langle\nabla_{\mathbf{w}_{n}}\mathcal{L}\left(\tilde{\mathbf{w}}_{n},\mathbf{0}\right),\overline{\mathbf{w}}_{n}-{\mathbf{w}}_{n}\right\rangle,$
where $\tilde{\mathbf{w}}_{n}=(1-t)\overline{\mathbf{w}}_{n}+t{\mathbf{w}}_{n}$, for some $t\in(0,1)$. Note that ${\mathbf{w}}_{n}-\overline{\mathbf{w}}_{n}=\left(\tilde{\mathbf{w}}_{n}-\overline{\mathbf{w}}_{n}\right)/t$. Thus,
$\mathcal{L}(\overline{\mathbf{w}}_{n},\mathbf{0})=\mathcal{L}(\mathbf{w}_{n},\mathbf{0})-\frac{1}{t}\left\langle\nabla_{\mathbf{w}_{n}}\mathcal{L}(\tilde{\mathbf{w}}_{n},\mathbf{0}),\tilde{\mathbf{w}}_{n}-\overline{\mathbf{w}}_{n}\right\rangle\leq\mathcal{L}(\mathbf{w}_{n},\mathbf{0}),$
where the inequality follows from (92), since $\|\tilde{\mathbf{w}}_{n}-\overline{\mathbf{w}}_{n}\|_{2}=t\|{\mathbf{w}}_{n}-\overline{\mathbf{w}}_{n}\|_{2}\leq\gamma$. Thus, $\overline{\mathbf{w}}_{n}$ is a local minimum. We next move towards proving the final part. Since $\overline{\mathbf{w}}_{n}$ is a local minimum of $\mathcal{L}({\mathbf{w}}_{n},\mathbf{0})$, there exists some $\gamma_{1}>0$ such that
$\mathcal{L}(\overline{\mathbf{w}}_{n},\mathbf{0})\leq\mathcal{L}({\mathbf{w}}_{n},\mathbf{0}),\text{ if }\|\mathbf{w}_{n}-\overline{\mathbf{w}}_{n}\|_{2}\leq\gamma_{1}.$
Further, since $\nabla^{2}_{\mathbf{w}_{n}}\mathcal{L}(\overline{\mathbf{w}}_{n},\mathbf{0})\succ\mathbf{0}$ and $\nabla^{2}_{\mathbf{w}_{n}}\mathcal{L}({\mathbf{w}}_{n},\mathbf{0})$ is continuous, there exists some $\gamma_{2}$ such that $\nabla^{2}_{\mathbf{w}_{n}}\mathcal{L}({\mathbf{w}}_{n},\mathbf{0})\succeq\mathbf{0},$ if $\|\mathbf{w}_{n}-\overline{\mathbf{w}}_{n}\|_{2}\leq\gamma_{2}$. Let $\gamma=\min(\gamma_{1},\gamma_{2})$, and choose any $\mathbf{w}_{n}$ such that $\|\mathbf{w}_{n}-\overline{\mathbf{w}}_{n}\|_{2}\leq\gamma$. By Taylor’s theorem we have
$\mathcal{L}(\overline{\mathbf{w}}_{n},\mathbf{0})=\mathcal{L}({\mathbf{w}}_{n},\mathbf{0})+\left\langle\nabla_{\mathbf{w}_{n}}\mathcal{L}({\mathbf{w}}_{n},\mathbf{0}),\overline{\mathbf{w}}_{n}-{\mathbf{w}}_{n}\right\rangle+\frac{1}{2}\left(\overline{\mathbf{w}}_{n}-{\mathbf{w}}_{n}\right)^{\top}\nabla^{2}_{\mathbf{w}_{n}}\mathcal{L}(\tilde{\mathbf{w}}_{n},\mathbf{0})\left(\overline{\mathbf{w}}_{n}-{\mathbf{w}}_{n}\right),$
where $\tilde{\mathbf{w}}_{n}=(1-t)\overline{\mathbf{w}}_{n}+t{\mathbf{w}}_{n}$, for some $t\in(0,1)$. Since $\|\tilde{\mathbf{w}}_{n}-\overline{\mathbf{w}}_{n}\|_{2}=t\|{\mathbf{w}}_{n}-\overline{\mathbf{w}}_{n}\|_{2}\leq\gamma$, we have $\nabla^{2}\mathcal{L}(\tilde{\mathbf{w}}_{n},\mathbf{0})\succeq\mathbf{0}$. Further, $\mathcal{L}(\overline{\mathbf{w}}_{n},\mathbf{0})\leq\mathcal{L}({\mathbf{w}}_{n},\mathbf{0})$, since $\|{\mathbf{w}}_{n}-\overline{\mathbf{w}}_{n}\|_{2}\leq\gamma\leq\gamma_{1}$. Thus,
$\left\langle\nabla\mathcal{L}({\mathbf{w}}_{n},\mathbf{0}),\overline{\mathbf{w}}_{n}-{\mathbf{w}}_{n}\right\rangle\leq 0,$
proving our claim. ∎
Appendix D Additional Lemma
In the following lemma we show that for two-homogeneous neural networks, the gradient flow dynamics of the NCF remains finite for any finite time.
Lemma D.1.
Let $\mathcal{H}(\mathbf{x};\mathbf{w})$ be a two-homogeneous neural network that satisfies first and third property of 1. Suppose $\mathbf{u}(t)$ is a solution of
$\frac{d\mathbf{u}}{dt}=\nabla\mathcal{N}_{\mathbf{z},\mathcal{H}}(\mathbf{u})=\mathcal{J}\left(\mathbf{X};\mathbf{u}\right)^{\top}\mathbf{z},\mathbf{u}(0)=\mathbf{u}_{0}.$ (93)
Then, for any finite time $T\geq 0$, $\|\mathbf{u}(T)\|_{2}$ is finite.
Proof.
Let $\beta=\sup\{\|\mathcal{H}(\mathbf{X};\mathbf{w})\|_{2}:\mathbf{w}\in\mathcal{S}^{k-1}\}$. Then
$\frac{1}{2}\frac{d{\|\mathbf{u}\|_{2}^{2}}}{dt}=\mathbf{u}^{\top}\dot{\mathbf{u}}=\mathbf{u}^{\top}\mathcal{J}\left(\mathbf{X};\mathbf{u}\right)^{\top}\mathbf{z}=2\mathcal{H}\left(\mathbf{X};\mathbf{u}\right)^{\top}\mathbf{z}\leq 2\beta\|\mathbf{u}\|_{2}^{2}\|\mathbf{z}\|_{2},$
where in the last equality we used Lemma B.1, and the inequality follows from Cauchy-Schwartz. Hence, for $T\geq 0$, we have
$\|\mathbf{u}(T)\|_{2}^{2}\leq\|\mathbf{u}_{0}\|_{2}^{2}+e^{4T\beta\|\mathbf{z}\|_{2}},$
proving our claim. ∎